### v1

# Quantum Advantage in Distributed Sensing with Noisy Quantum Networks

Abstract
We show that quantum advantage in distributed sensing can be achieved with noisy quantum networks. When using depolarized GHZ states as the probe, we derive a closed-form fidelity threshold to achieve advantage over the optimal local sensing strategy. The threshold indicates that while entanglement is needed for this quantum advantage, genuine multipartite entanglement is generally unnecessary. We further explore the impacts from imperfect local entanglement generation and local measurement constraint, and our results imply that the quantum advantage is more robust against quantum network imperfections than local operation errors. Finally, we demonstrate that the quantum advantage in distributed sensing can be achieved with a three-node quantum network using practical protocol stacks through simulations with SeQUeNCe, an open-source, customizable quantum network simulator.
preprint: APS/123-QED
Introduction.—Distributed quantum sensing (DQS)[1, 2, 3] is one of the most important applications of quantum networks [4, 5]. It is expected to surpass classical sensing techniques in areas ranging from magnetometry [6, 7, 8, 9], phase imaging [10], precision clocks [11], energy applications [12], all the way to the exploration of fundamental physics [13, 14], including search for dark matter [15] and measuring stability of fundamental constants [16].
Over the past decade, DQS has attracted great efforts on the derivation of its ultimate limit [17, 18, 19, 20, 21, 22]. Various DQS protocols have also been proposed under ideal conditions [23, 24, 25, 26]. On the other hand, as the principle of DQS has been experimentally demonstrated in small-scale matter-based quantum systems [27], and the experimental implementation of entanglement distribution networks is quickly advancing [28, 29, 30], it is crucial to analyze the feasibility of DQS with realistic, noisy quantum networks. Despite its importance and necessity, the analysis of DQS with imperfection, especially the state preparation error that is unavoidable in realistic quantum networks, has only emerged in recent years and is insufficient. For instance, Sekatski et al. [31] and some later works [32, 33, 34] considered DQS with noise in signals, rather than the unavoidable state preparation error over a realistic quantum network. Fadel et al. [35] studied partitioning atomic ensembles for DQS, but again did not consider state preparation noise, and the assumed physical scenario is still small-scale. Van Milligen et al. [36] considered a quantum network scenario with focus on the probabilistic nature of network operations.
In this work, we focus on practical DQS scenarios and study the possibility of realizing quantum advantage in DQS with noisy quantum networks. We first specify the initial probe state, the parameter encoding process, and the estimation objective. We then derive closed-form upper bounds for DQS estimation accuracy and theoretically demonstrate the possibility of quantum advantage over the optimal local sensing protocol. The impacts from imperfect local entanglement generation when scaling up the number of local sensors, and local measurement constraint are also analyzed. In addition, we perform a quantum network simulation which shows that DQS quantum advantage can in principle be achieved by realistic quantum network configurations.
Problem formulation.—We consider estimating the average of spatially distributed parameters [37]. This task can be decomposed into three steps: (i) preparation of the initial probe state, (ii) encoding the parameters on the probe state, and (iii) performing measurement on the encoded state to extract information about the parameters. We assume unitary parameter encoding in the standard phase accumulation form: $U(x)=\exp\left[-i\left(\sum_{i=0}^{d-1}x_{i}H_{i}\right)\right]$, where $x=(x_{1},\dots,x_{d})^{T}\in\mathbb{R}^{d}$ is an array of $d$ parameters located at $d$ sensor nodes in the network, and $H_{i}=\frac{1}{2}\sum_{k=0}^{n-1}\sigma^{(i,k)}_{z}$ is the z-component of collective spin for the $n$ local qubit sensors on node $i$. Although in practice there will always be decoherence during quantum dynamics, in a DQS setup the time scale for quantum network operations could be significantly longer than local quantum operations, and therefore the local parameter encoding process could be much less noisy than the initial probe state preparation over the quantum network. By assuming a noiseless encoding process, it is also easier to study the impact of network imperfections and distinguish it from the effect of the encoding process on the DQS performance.
Under the aforementioned unitary encoding channel, the optimal probe state for this problem has been shown to be the global GHZ state  [2, 1]. Therefore, we choose this global GHZ state as the target initial probe state to prepare in the quantum network of sensors. The preparation of the probe state over a quantum network can be decomposed into two steps. Firstly the quantum network will distribute a $d$-qubit GHZ state across all $d$ sensor nodes, with each node having one qubit. Each node can then perform local entanglement generation [38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48] to extend the size of the global GHZ state: By entangling $(n-1)$ additional quantum sensors per node, the global GHZ state will have $nd$ qubits in total. Given noisy quantum networks, we consider that the prepared initial probe state is mixed. Motivated by Pauli twirling [49, 50, 51] which is able to transform error channels into Pauli channels, and stabilizer twirling [52] which can cancel off-diagonal density matrix elements under a stabilizer basis, we assume that the initial probe state is GHZ-diagonal: $\rho_{0}=\sum_{a}\lambda_{a}|\psi_{a0}\rangle\langle\psi_{a0}|$ with all $\lambda_{a}$ being non-zero, where $a$ is the index of the pure GHZ basis state $|\psi_{a0}\rangle$. This Pauli error model has also been utilized in a recent study of single-parameter quantum metrology with graph states [53]. According to the assumed GHZ-diagonal form of initial probe state, we can analytically derive the quantum Fisher information (QFI) [54] which characterizes the lower bound of parameter estimation variance, the quantum Cramér-Rao bound (QCRB).
Quantum Fisher information.—To derive the QFI for the average of all local parameters, we start with the QFI matrix (QFIM) [55, 56, 57, 58] for the $d$ local parameters. According to the assumed parametrization unitary, for a general GHZ-diagonal $nd$-qubit initial probe state $\rho_{0}$ where each of the $d$ sensor nodes holds $n$ qubit sensors, the QFIM is always proportional to a matrix of ones:
$\mathcal{F}(x)=(1-C)n^{2}\begin{pmatrix}1&1&\dots&1\\ 1&1&\dots&1\\ \vdots&\vdots&\ddots&\vdots\\ 1&1&\dots&1\\ \end{pmatrix},$ (1)
where $\mathcal{F}(x)$ denotes the QFIM for parameters $x$, and $C$ captures the quality of the initial probe state, and it is analytically calculated as:
$C=\sum_{(a,b)\in\mathcal{S}}\frac{2\lambda_{a0}\lambda_{b0}}{\lambda_{a0}+\lambda_{b0}},$ (2)
where $\lambda_{a0}$ is the eigenvalue corresponding to GHZ basis state $|\psi_{a0}\rangle$, and $\mathcal{S}$ denotes the set of index pairs $(a,b)$ with double counting, such that $|\psi_{a0}\rangle$ and $|\psi_{b0}\rangle$ are GHZ states expressed as superpositions of the same pair of computational basis states but with opposite relative phase. For instance, for 3-qubit GHZ states $(|000\rangle\pm|111\rangle)/\sqrt{2}$ is such a pair. Note that $C$ is not necessarily a constant and in general depends on $n$, and its dependence is determined by error models.
Our parameter to estimate is $\theta_{1}=v_{1}^{T}x$ where $v\propto(1,\dots,1)^{T}$. For concreteness, we follow the convention in [2, 1] and choose $v_{1}=(1,\dots,1)/\sqrt{d}$ which is normalized under the 2-norm. Then we are able to transform the QFIM for the new global parameter $\theta_{1}$, where the transformation could be operationally realized by constructing additional $(d-1)$ normalized vectors $v_{2},\dots,v_{d}$, s.t. $v_{i}^{T}v_{j}=\delta_{ij}$ [55]. Consequently, we have:
$\mathcal{F}(\theta)=d(1-C)n^{2}\begin{pmatrix}1&0&\dots&0\\ 0&0&\dots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\dots&0\\ \end{pmatrix},$ (3)
where $\theta=(v_{1}^{T}x,\dots,v_{d}^{T}x)$, and the only non-zero entry is the QFI for $\theta_{1}$ of our interest: $\mathcal{F}(\theta_{1})=d(1-C)n^{2}$. The detailed derivation for the above results can be found in Sec. II of the Supplemental Material [59].
Condition for quantum advantage.—To demonstrate quantum advantage, the comparison baseline should be the optimal local sensing strategy where each sensor node can estimate the local parameter to the best accuracy possible under the resource constraints, and we then use the estimated values to approximate their average.
It is well known that the GHZ state is the optimal probe state for local phase estimation [60], and indeed using separate GHZ states on each sensor node is the best local strategy (without any quantum communication between sensor nodes) for estimating the average of local parameters [2, 1]. The variance of this indirect estimation can be calculated through propagation of error [57]: $\mathrm{Var}_{\mathrm{local}}(\hat{\theta}_{1})=\sum_{l=1}^{d}\left(\frac{\partial\theta_{1}}{\partial x_{l}}\right)^{2}\mathrm{Var}_{\mathrm{local}}(\hat{x}_{l})$, where we use the QCRB of local GHZ state with $n$ qubits so that $\mathrm{Var}_{\mathrm{local}}(\hat{x}_{l})=1/(n^{2}\mu)$, where $\mu$ is the number of samples. Noticing that $\partial\theta_{1}/\partial x_{l}=1/\sqrt{d}$, we immediately have $\mathrm{Var}_{\mathrm{local}}(\hat{\theta}_{1})=1/(n^{2}\mu)$. Therefore, a quantum advantage of entangling sensor nodes with quantum networks is achievable when:
$\eta=d(1-C)>1.$ (4)
Next, we will evaluate $\eta$ for $nd$-qubit GHZ state created from the initial noisy $d$-qubit GHZ state with noiseless and noisy local entanglement generation, respectively.
Noiseless local entanglement generation.—We first consider noiseless local entanglement generation, which can be interpreted as applying perfect CNOT gates between the qubit sensor initially entangled with other sensor nodes and other local qubit sensors prepared in $|0\rangle$ state to entangle, with the former being the control. This assumption allows us to further decouple the network imperfections and local errors, and thus to understand the limit on the amount of network imperfections beyond which there cannot be any quantum advantage for DQS. Without loss of generality, we assume that the initial noisy $d$-qubit GHZ state is full-rank, and in the form of a depolarized GHZ state, i.e. a mixture of pure GHZ state and maximally mixed state, $\rho_{\mathrm{dp}}(F)$, where $F$ is the fidelity to the pure GHZ state. We can then derive a closed-form fidelity threshold, below which there will be no quantum advantage in DQS:
$F_{\mathrm{th,dp}}=2^{-d}+\frac{(2^{d}-1)\left(2^{d}-2+\sqrt{(2^{d}-2)^{2}+2^{d+3}d}\right)}{2^{2d+1}d}.$ (5)
We see that in the asymptotic regime of large $d$, the above fidelity threshold reduces to $1/d$. Moreover, the quantum advantage in DQS has close relationship with quantum entanglement, and the entanglement properties of the depolarized GHZ state have been well studied: $d$-qubit depolarized GHZ states are not completely separable if $F>3/(2^{d}+2)$ [61], and genuine multipartite entangled (GME) if $F>1/2$ [62]. It can be shown that $F_{\mathrm{th,dp}}>3/(2^{d}+2)$ for $d\geq 2$, which means that entanglement is necessary for DQS quantum advantage. Meanwhile, except for the special case of $\left.F_{\mathrm{th,dp}}\right|_{d=3}\approx 0.50963>1/2$, $F_{\mathrm{th,dp}}<1/2$ for $d>3$, which suggests that GME is in principle unnecessary to demonstrate quantum advantage in DQS. Further details can be found in Sec. III.A of the Supplemental Material [59].
Figure 1: Visualization of $\eta_{\mathrm{dp}}$ as function of the number of local qubits $n$, for some choices of initial fidelity $F$ and local entanglement generation quality $k$. We fix $d=3$ while varying $F$ and $k$. Dashed curves correspond to $F=0.8$; solid curves correspond to $F=0.9$; dot-dashed curves correspond to $F=0.99$. Blue color represents $k=0.9999$; yellow color represents $k=0.999$; green color represents $k=0.99$.
Imperfect local entanglement generation.—We now start to include imperfections in local entanglement generation. Specifically, we consider the following phenomenological model to reflect the fidelity decrease when adding more qubits to the GHZ state: We assume that the state after local entanglement generation is an $nd$-qubit depolarized GHZ state, while the fidelity is modified according to the number of local qubits as $F(n)=k^{n-1}F$, where $F$ is the fidelity of the initial $d$-qubit GHZ state, and $k\in(0,1)$ is a parameter which describes the quality of local entanglement generation and thus the larger the better. The intuition behind this phenomenological model is the usage of noisy CNOT gates to generate GHZ states, and $k$ can be interpreted as gate fidelity. We can then express $C$ in Eqn. 2 as:
$C_{\mathrm{dp}}(F,d,n,k)=\frac{(1-k^{n-1}F)(4^{nd}k^{n-1}F-2^{nd}-2)}{[(2^{nd}-2)k^{n-1}F+1](2^{nd}-1)}.$ (6)
It can be shown that, as intuitively expected, $C_{\mathrm{dp}}$ decreases monotonically as initial fidelity $F$, number of local parameters $d$, and local entanglement generation quality $k$, increase for all $n\geq 1$, if $F,k>2^{-d}$. On the other hand, the dependence of $C_{\mathrm{dp}}$ on number of local quantum sensors $n$ is more complicated. We visualize $\eta_{\mathrm{dp}}=d(1-C_{\mathrm{dp}})$ with varying $n$ under different parameter choices in Fig. 1.
From the plot we can verify that the relative advantage $\eta_{\mathrm{dp}}$ is larger when $F,k$ increase. We can also see non-monotonic behavior of $\eta_{\mathrm{dp}}$ when $k$ is sufficiently high in comparison to $F$, which indicates an increase in relative advantage for small local sensor numbers. The intersection point at which $\eta_{\mathrm{dp}}=1$ determines the maximal number of local quantum sensors $n_{\mathrm{max}}$ for quantum advantage over the optimal local sensing strategy to be potentially demonstrated. We observe that $n_{\mathrm{max}}$ does not change much when varying $F$ given fixed $k$, but it changes significantly when fixing $F$ and varying $k$, which suggests that the initial fidelity $F$ has less impact on $n_{\mathrm{max}}$ than the local entanglement generation quality $k$. This can be understood by considering the fidelity threshold for the $nd$-qubit depolarized GHZ state to be advantageous over the optimal local strategy, which also scales as $1/d$ and is almost independent of $n$. Imperfect local entanglement generation will decrease the GHZ fidelity to the threshold at $n_{\mathrm{max}}$. Therefore, we expect $k^{n_{\mathrm{max}}-1}F\approx 1/d$, which gives $n_{\mathrm{max}}\approx-\ln(dF)/\ln(k)$. This also implies robustness of the quantum advantage in DQS against the imperfection of probe state generation over realitic quantum networks. See more details in Sec. III.B of the Supplemental Material [59].
Local measurement constraint.—It is possible that the optimal measurement to saturate the QCRB is entangling between the sensor nodes. In the DQS setup, entangling measurement needs additional remote entanglement as a resource to implement. Therefore, we want to use only local operations and classical communication (LOCC) [63] to extract information in DQS problems. However, in general LOCC is not guaranteed to saturate the QCRB in DQS [64]. Here we explicitly consider the local measurement constraint.
It is known that $M=\sigma_{x}^{\otimes nd}$ is the optimal observable for $nd$-qubit pure GHZ probe state under unitary z-direction phase accumulation [60], and in principle each sensor node only needs to perform measurement of the local observable $\sigma_{x}^{\otimes n}$. However, this optimal measurement under ideal conditions is genuinely useless whenever there is any noise in the GHZ state as probe. Specifically, the variance of parameter estimation will diverge when taking the limit of small local parameters. We note that this conclusion is also implied by numerical results in another recent work by Cao and Wu [65]. In Sec. IV.A of the Supplemental Material [59] we analytically demonstrate this property for the depolarized GHZ state, while it holds generally for any GHZ-diagonal state with non-unit fidelity.
Motivated by our problem formulation that the parameter to estimate is encoded through z-direction phase accumulation, we further explore the optimization over a subset of local measurements, i.e. the tensor product of single-qubit measurements along a direction on the equator of the Bloch sphere: $M(\alpha)=\left[O(\alpha)\right]^{\otimes nd}$ where $O(\alpha)=|\psi^{+}(\alpha)\rangle\langle\psi^{+}(\alpha)|-|\psi^{-}(\alpha)\rangle\langle\psi^{-}(\alpha)|$ with $|\psi^{\pm}(\alpha)\rangle=(|0\rangle\pm e^{i\alpha}|1\rangle)$ (thus $O(\alpha)=e^{i\alpha}|1\rangle\langle 0|+e^{-i\alpha}|0\rangle\langle 1|$), characterized by the azimuthal angle $\alpha$. For an $nd$-qubit depolarized GHZ state, the optimal azimuthal angle is $\alpha_{\mathrm{opt}}=\frac{2l+1}{2nd}\pi,\ l\in\mathbb{Z}$. Note that the optimal azimuthal angle depends on number of local quantum sensors $n$ and sensor node number $d$. Moreover, the estimation variance diverges quickly when $\alpha$ deviates from $\alpha_{\mathrm{opt}}$. This implies that the accuracy of local operation is extremely important, and potentially more important than the quality of the entangled states distributed by quantum networks.
Figure 2: Visualization of fidelity thresholds. The blue solid curve denotes the fidelity threshold given by the QFI, and the threshold given by optimized local azimuthal measurement is shown as the red solid curve. The horizontal and curved black dashed lines again represent the fidelity thresholds for $d$-qubit depolarized GHZ state to be GME and non-fully separable, respectively.
The fidelity threshold for an $nd$-qubit depolarized GHZ state to be advantageous over the optimal local strategy when using the optimized azimuthal measurement is:
$F_{\mathrm{th},M(\alpha_{\mathrm{opt}})}(n)=\frac{2^{nd}+\sqrt{d}-1}{2^{nd}\sqrt{d}}.$ (7)
The case of $n=1$ for Eq. 7 corresponds to noiseless local entanglement generation, similar to Eq. 5. We thus visualize and compare both fidelity thresholds in Fig. 2. The fidelity threshold in Eq. 7 is always higher than that given by Eq. 5, suggesting that the optimized local azimuthal measurement does not saturate the QCRB. On the other hand, although for small problem sizes $d<5$ the initial GHZ state needs to be GME to demonstrate quantum advantage if we use the optimized azimuthal measurement, when the problem size grows the requirement of initial fidelity drops, and in general the initial $d$-qubit GHZ state does not have to be GME. More details on the azimuthal measurement optimization can be found in Sec. IV.B of the Supplemental Material [59].
Table 1: Simulation results for the three scenarios.
$p$ $\eta$ $\tilde{\eta}$ $F$
Scenario 1 $\approx 0.02$ $\approx 0.95<1$ $\approx 0.02<1$ 0.591
Scenario 2 $\approx 0.72$ $\approx 1.57>1$ $\approx 1.13>1$ 0.732
Scenario 3 $\approx 0.81$ $\approx 2.19>1$ $\approx 1.77>1$ 0.854
Quantum network simulation.—Following the above analytics, we simulate GHZ state distribution given realistic first-generation quantum repeater network [66, 67, 68] protocol stacks with an open-source, customizable quantum network simulator, SeQUeNCe [69]. We consider a simple 3-node network topology, where a center node directly connects the other two end nodes through optical fibers, with a Bell state measurement station in the middle of each fiber link. First, bipartite entanglement links are estabilished between the center node and the other nodes. Then LOCC, such as gate teleportation [70, 71, 72] and graph state fusion [73, 74], are able to generate a GHZ state across all network nodes.
In our simulations we consider three scenarios, and thus three sets of system parameter values: $\{0.01\mathrm{s},0.1\mathrm{s},1\mathrm{s}\}$ for memory coherence times, $\{0.05,0.1,0.5\}$ for memory efficiency, and $\{0.8,0.85,0.9\}$ for raw entanglement fidelity. The first simulation scenario uses the first value in each of the above sets, and so on. Meanwhile, we fix other system parameters, especially: a 1s time interval allowed for entanglement distribution, a 10km distance between the center node and end nodes, 10 memories per end node and 20 memories on the center node so that entanglement purification [75, 76] is possible. We emphasize that the parameter values are not chosen from a specific reported experiment, but are selected according to the general vision of the state of the art and the potential future development of various candidate platforms for quantum networking, including solid state systems [28, 30, 77], atomic systems [29, 78], and superconducting systems [79]. Details of our quantum network simulation can be found in Sec. VII of the Supplemental Material [59].
We characterize the performance of probe state distribution by three figures of merit, namely the success probability of distributing the 3-qubit GHZ state, $p$, the relative advantage, $\eta$, and the normalized relative advantage, $\tilde{\eta}=p\cdot\eta$, which takes into account failure in entanglement distribution (See Sec. V of the Supplemental Material [59]). For each scenario we repeat the simulation 1000 times, and $\eta$ is calculated based on the average of density matrices of successfully distributed GHZ states under ensemble interpretation. The results are collected in Table 1, from which it is clear that the most modest parameter choice does not permit quantum advantage, while when hardware performance improves quantum advantage becomes possible without changing the realistic quantum network protocol stack. We also note the average fidelity $F$ of the successfully generated states for each scenario; as expected, the fidelity increases as the network parameters improve.
Conclusion and discussion.—In summary, we perform extensive analytical studies on the impact of imperfect GHZ state distributed by noisy quantum networks on DQS. Our results offer new insights into realistic DQS, and reveal the relation between entanglement and DQS quantum advantage. We also simulate the GHZ state distribution process over a 3-node quantum network, demonstrating the possibility of DQS quantum advantage with realistic quantum network stacks. The new features we develop in SeQUeNCe to reflect imperfections in entanglement distribution are completely open-source [80], and can thus serve as valuable resources for future quantum network research. We leave more detailed quantum network simulation as a followup work.
We have primarily focused on probe state preparation errors, while the parameter encoding dynamics can also be noisy due to sensor decoherence. It could be interesting to evaluate the impact of sensor decoherence after state preparation [81], where the time dependence of estimation accuracy gain [82, 83, 84] becomes important. We note that quantum error correction [85, 86, 87, 88] offers a good opportunity in fighting against noise in sensing process. Meanwhile, protocols such as continuous entanglement distribution [89, 90, 91] might be necessary to reduce latency of probe state preparation over quantum networks. In addition, optimization of Bell-state-based graph state distribution [92, 93, 94, 95, 96, 97, 98, 99, 100] is important for larger scale DQS. Finally, although we mainly assume finite-dimensional matter-based quantum sensors, photonic systems [101, 102, 103, 104, 105, 106] are also playing an important role in DQS.
Acknowledgments.—A.Z. would like to thank Tian-Xing Zheng and Boxuan Zhou for helpful discussions. This material is based upon work supported by the U.S. Department of Energy Office of Science National Quantum Information Science Research Centers. Work performed at the Center for Nanoscale Materials, a U.S. Department of Energy Office of Science User Facility, was supported by the U.S. DOE, Office of Basic Energy Sciences, under Contract No. DE-AC02-06CH11357. A.Z. and T.Z. are also supported by the NSF Quantum Leap Challenge Institute for Hybrid Quantum Architectures and Networks (NSF Grant No. 2016136), and the Marshall and Arlene Bennett Family Research Program.