### v3

# Quantum Advantage in Distributed Sensing with Noisy Quantum Networks

Abstract
It is critically important to analyze the achievability of quantum advantage under realistic imperfections. In this work, we show that quantum advantage in distributed sensing can be achieved with noisy quantum networks which can only distribute noisy entangled states. We derive a closed-form expression of the quantum Fisher information (QFI) for estimating the average of local parameters using GHZ-diagonal probe states, an important distributed sensing prototype. From the QFI we obtain the necessary condition to achieve quantum advantage over the optimal local sensing strategy, which can also serve as an optimization-free entanglement detection criterion for multipartite states. In addition, we prove that genuine multipartite entanglement is neither necessary nor sufficient through explicit examples of depolarized and dephased GHZ states. We further explore the impacts from imperfect local entanglement generation and local measurement constraint, and our results imply that the quantum advantage is more robust against quantum network imperfections than local operation errors. Finally, we demonstrate that the probe state with potential for quantum advantage in distributed sensing can be prepared by a three-node quantum network using practical protocol stacks through simulations with SeQUeNCe, an open-source, customizable quantum network simulator.
preprint: APS/123-QED
Introduction.—Distributed quantum sensing (DQS) [1, 2, 3] is one of the most important applications of quantum networks [4, 5]. It is expected to surpass classical sensing techniques in areas ranging from magnetometry [6, 7, 8, 9], phase imaging [10], precision clocks [11], energy applications [12], all the way to the exploration of fundamental physics [13, 14], including search for dark matter [15] and measuring stability of fundamental constants [16]. Over the past decade, DQS has attracted great theoretical interests [17, 18, 19, 20, 21, 22, 23, 24]. Various DQS protocols have been proposed under ideal conditions [25, 26, 27, 28, 29, 30, 31]. On the other hand, along with the experimental demonstration of concept in small-scale matter systems [32], the analysis of realistic DQS has also emerged recently, for instance, DQS with noise in signals [33, 34, 35, 36], atomic ensemble partition for DQS [37], and DQS with probabilistic quantum network operations [38]. However, the state-of-the-art progress in quantum networking [39, 40, 41] demonstrates that remote entanglement has much higher infidelity than local operations, so the process of initial state preparation will be the major error source for DQS in the near term. Despite its critical importance and necessity, the systematic analysis of DQS with imperfect state preparation surprisingly remains largely unexplored.
In this work, we focus on studying the possibility of realizing quantum advantage in DQS with noisy quantum networks which can only distribute noisy entangled states. We first specify the initial probe state, the parameter encoding process, and the estimation objective. We then derive the condition for quantum advantage over the optimal local sensing protocol, and demonstrate with the example of depolarized GHZ state as the initial probe. The impacts from imperfect local entanglement generation when scaling up the number of local sensors, and local measurement constraint are also analyzed. In addition, we perform a quantum network simulation which shows that DQS quantum advantage can in principle be achieved by realistic quantum network configurations.
Problem formulation.—We consider estimating the average of spatially distributed parameters [42]. This task can be decomposed into three steps: (i) preparation of the initial probe state, (ii) encoding the parameters on the probe state, and (iii) performing measurement on the encoded state to extract information about the parameters. We assume unitary parameter encoding in the standard phase accumulation form: $U(x)=\exp\left[-i\left(\sum_{i=0}^{d-1}x_{i}H_{i}\right)\right]$, where $x=(x_{1},\dots,x_{d})^{T}\in\mathbb{R}^{d}$ is an array of $d$ parameters located at $d$ sensor nodes in the network, and $H_{i}=\frac{1}{2}\sum_{k=0}^{n-1}\sigma^{(i,k)}_{z}$ is the z-component of collective spin for the $n$ local qubit sensors on node $i$. Although in practice there will always be decoherence during quantum dynamics, in a DQS setup the time scale for quantum network operations could be significantly longer than local quantum operations, and therefore the local parameter encoding process could be much less noisy than the initial probe state preparation over the quantum network. By assuming a noiseless encoding process, it is also easier to study the impact of network imperfections and distinguish it from the effect of the encoding process on the DQS performance.
Under the aforementioned unitary encoding channel, the optimal probe state for this problem has been shown to be the global GHZ state  [2, 1], i.e., a GHZ state $|\mathrm{GHZ}\rangle=(|0\dots 0\rangle+|1\dots 1\rangle)/\sqrt{2}$ involving all the sensors on each sensor node. Therefore, we choose this global GHZ state as the target initial probe state to prepare in the quantum network of sensors. The preparation of the probe state over a quantum network can be decomposed into two steps. Firstly, the quantum network will distribute a $d$-qubit GHZ state across all $d$ sensor nodes, with each node having one qubit. Each node can then perform local entanglement generation [43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53] to extend the size of the global GHZ state: By entangling $(n-1)$ additional quantum sensors per node, the global GHZ state will have $nd$ qubits in total. Note that only $(d-1)$ Bell pairs are needed to distribute the $d$-qubit GHZ state over the $d$ nodes (see Sec. VI of the Supplemental Material [54]). Therefore, such global GHZ states are easy to generate by quantum networks, which further justifies why we choose them as the probe states. Given noisy quantum networks, we consider that the prepared initial probe state is mixed. Motivated by Pauli twirling [55, 56, 57] which is able to transform error channels into Pauli channels, and stabilizer twirling [58] which can cancel off-diagonal density matrix elements under a stabilizer basis, we assume that the initial probe state is GHZ-diagonal: $\rho_{0}=\sum_{a}\lambda_{a}|\psi_{a0}\rangle\langle\psi_{a0}|$ with all $\lambda_{a}$ being non-zero, where $a$ is the index of the pure GHZ basis state $|\psi_{a0}\rangle$. This Pauli error model has also been utilized in a recent study of single-parameter quantum metrology with graph states [59]. According to the assumed GHZ-diagonal form of initial probe state, we can analytically derive the quantum Fisher information (QFI) [60] which characterizes the lower bound of parameter estimation variance, the quantum Cramér-Rao bound (QCRB).
Quantum Fisher information.—To derive the QFI for the average of all local parameters, we start with the QFI matrix (QFIM) [61, 62, 63, 64] for the $d$ local parameters. According to the assumed parametrization unitary, for a general GHZ-diagonal $nd$-qubit initial probe state $\rho_{0}$ where each of the $d$ sensor nodes holds $n$ qubit sensors, the QFIM for parameters $x$ is $\mathcal{F}(x)=(1-C)n^{2}J_{d}$, where $J_{d}$ is the $d$-by-$d$ matrix of ones, and $C$ captures the quality of the initial probe state, and it is analytically calculated as:
$C=\sum_{(a,b)\in\mathcal{S}}\frac{2\lambda_{a0}\lambda_{b0}}{\lambda_{a0}+\lambda_{b0}},$ (1)
where $\lambda_{a0}$ is the eigenvalue corresponding to GHZ basis state $|\psi_{a0}\rangle$, and $\mathcal{S}$ denotes the set of index pairs $(a,b)$ with double counting, such that $|\psi_{a0}\rangle$ and $|\psi_{b0}\rangle$ are GHZ states expressed as superpositions of the same pair of computational basis states but with opposite relative phase. For instance, for 3-qubit GHZ states $(|000\rangle\pm|111\rangle)/\sqrt{2}$ is such a pair. Note that $C$ is not necessarily a constant and in general depends on $n$, and its dependence is determined by error models.
Our parameter to estimate is $\theta_{1}=v_{1}^{T}x$ where $v\to(1,\dots,1)^{T}$. For concreteness, we follow the convention in [2, 1] and choose $v_{1}=(1,\dots,1)^{T}/\sqrt{d}$ which is normalized under the 2-norm. Then we are able to transform the QFIM for the new global parameter $\theta_{1}$, where the transformation could be operationally realized by constructing additional $(d-1)$ normalized vectors $v_{2},\dots,v_{d}$, s.t. $v_{i}^{T}v_{j}=\delta_{ij}$ [61]. Consequently, we have:
$\mathcal{F}(\theta)=d(1-C)n^{2}\begin{pmatrix}1&0&\dots&0\\ 0&0&\dots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\dots&0\\ \end{pmatrix},$ (2)
where $\theta=(v_{1}^{T}x,\dots,v_{d}^{T}x)$, and the only non-zero entry is the QFI for $\theta_{1}$ of our interest: $\mathcal{F}(\theta_{1})=d(1-C)n^{2}$, which incorporates effects from both the network ($d$) and local resources ($n^{2}$). The details can be found in Sec. II of the Supplemental Material [54].
Condition for quantum advantage.—To demonstrate quantum advantage, the comparison baseline should be the optimal local sensing strategy where each sensor node can estimate the local parameter to the best accuracy possible under the resource constraints, and we then use the estimated values to approximate their average.
It is well known that the GHZ state is the optimal probe state for local phase estimation [65], and indeed using separate GHZ states on each sensor node is the best local strategy (without any quantum communication between sensor nodes) for estimating the average of local parameters [2, 1]. The variance of this indirect estimation can be calculated through propagation of error [63]: $\mathrm{Var}_{\mathrm{local}}(\hat{\theta}_{1})=\sum_{l=1}^{d}\left(\frac{\partial\theta_{1}}{\partial x_{l}}\right)^{2}\mathrm{Var}_{\mathrm{local}}(\hat{x}_{l})$, where we use the QCRB of local GHZ state with $n$ qubits so that $\mathrm{Var}_{\mathrm{local}}(\hat{x}_{l})=1/(n^{2}\mu)$, where $\mu$ is the number of samples. Noticing that $\partial\theta_{1}/\partial x_{l}=1/\sqrt{d}$, we immediately have $\mathrm{Var}_{\mathrm{local}}(\hat{\theta}_{1})=1/(n^{2}\mu)$. Therefore, on top of the scaling advantage from increasing local resources, the additional relative quantum advantage of entangling sensor nodes with quantum networks is achievable when:
$\eta\equiv d(1-C)>1.$ (3)
Since entanglement in the initial probe state across sensor nodes is necessary for DQS quantum advantage, the above condition can also be interpreted as an operationally meaningful, optimization-free entanglement detection criterion for arbitrary $d$-qubit states. See details in Sec. II of the Supplemental Material [54]. Next, we will evaluate $\eta$ for $nd$-qubit GHZ state created from the initial noisy $d$-qubit GHZ state with noiseless and noisy local entanglement generation, respectively.
Noiseless local entanglement generation.—We first consider noiseless local entanglement generation, which can be interpreted as applying perfect CNOT gates between the qubit sensor initially entangled with other sensor nodes and other local qubit sensors prepared in $|0\rangle$ state to entangle, with the former being the control. This assumption results in that the $nd$-qubit GHZ state has the same fidelity as the $d$-qubit GHZ state. We are then allowed to decouple the network imperfections and local errors, and thus to understand the limit on the amount of network imperfections beyond which there cannot be any quantum advantage for DQS. While the analysis of $\eta$ is applicable to arbitrary GHZ-diagonal states, here for concreteness we assume a canonical full-rank error model for the initial $d$-qubit GHZ state, a depolarized GHZ state as a mixture of pure GHZ state and maximally mixed state, $\rho_{\mathrm{dp}}(F)$, where $F$ is the fidelity to the pure GHZ state. We can then substitute the $C$ for $d$-qubit depolarized GHZ state in Eq. 3 and let $\eta=1$ to derive a closed-form fidelity threshold, as a requirement on the quantum network performance for quantum advantage in DQS:
$F_{\mathrm{th,dp}}=2^{-d}+\frac{(2^{d}-1)\left(2^{d}-2+\sqrt{(2^{d}-2)^{2}+2^{d+3}d}\right)}{2^{2d+1}d}.$ (4)
We see that in the asymptotic regime of large $d$, the above fidelity threshold reduces to $1/d$. Moreover, the quantum advantage in DQS has a close relationship with quantum entanglement, and the entanglement properties of the depolarized GHZ state have been well studied: $d$-qubit depolarized GHZ states are not completely separable if $F>3/(2^{d}+2)$ [66], and genuine multipartite entangled (GME) if $F>1/2$ [67]. It can be shown that $F_{\mathrm{th,dp}}>3/(2^{d}+2)$ for $d\geq 2$, which means that entanglement is necessary for DQS quantum advantage. Meanwhile, except for the special case of $\left.F_{\mathrm{th,dp}}\right|_{d=3}\approx 0.50963>1/2$, $F_{\mathrm{th,dp}}<1/2$ for $d>3$, which suggests that GME is in principle unnecessary to demonstrate quantum advantage in DQS. Moreover, via analysis of rank-2 dephased GHZ states, we can show that GME is not sufficient for DQS quantum advantage either. Details can be found in Sec. III.A of the Supplemental Material [54].
Figure 1: Visualization of $\eta_{\mathrm{dp}}$ as function of the number of local qubits $n$, for different initial fidelity $F$ and local entanglement generation quality $k$, with fixed $d=3$.
Imperfect local entanglement generation.—We now start to include imperfections in local entanglement generation. Specifically, we consider the following phenomenological model to reflect the fidelity decrease when adding more qubits to the GHZ state: We assume that the state after local entanglement generation is an $nd$-qubit depolarized GHZ state, while the fidelity is modified according to the number of local qubits as $F(n)=k^{n-1}F$, where $F$ is the fidelity of the initial $d$-qubit GHZ state, and $k\in(0,1)$ is a parameter which describes the quality of local entanglement generation and thus the larger the better. The intuition behind this phenomenological model is the usage of noisy CNOT gates to generate GHZ states, and $k$ can be interpreted as gate fidelity. We can again evaluate $C$ in Eq. 1 as a function of the four system parameters $C_{\mathrm{dp}}(F,d,n,k)$, whose expression can be found in Sec. III.B of the Supplemental Material [54]. It can be shown that, as intuitively expected, $C_{\mathrm{dp}}$ decreases monotonically as initial fidelity $F$, number of local parameters $d$, and local entanglement generation quality $k$, increase for all $n\geq 1$, if $F,k>2^{-d}$. On the other hand, the dependence of $C_{\mathrm{dp}}$ on number of local quantum sensors $n$ is more complicated. We visualize $\eta_{\mathrm{dp}}=d(1-C_{\mathrm{dp}})$ with varying $n$ under different parameter choices in Fig. 1.
From the plot we can verify that the relative advantage $\eta_{\mathrm{dp}}$ is larger when $F,k$ increase. We can also see non-monotonic behavior of $\eta_{\mathrm{dp}}$ when $k$ is sufficiently high in comparison to $F$, which indicates an increase in relative advantage for small local sensor numbers. The intersection point at which $\eta_{\mathrm{dp}}=1$ determines the maximal number of local quantum sensors $n_{\mathrm{max}}$ for quantum advantage over the optimal local sensing strategy to be potentially demonstrated. We observe that $n_{\mathrm{max}}$ does not change much when varying $F$ given fixed $k$, but it changes significantly when fixing $F$ and varying $k$, which suggests that the initial fidelity $F$ has less impact on $n_{\mathrm{max}}$ than the local entanglement generation quality $k$. This can be understood by considering the fidelity threshold for the $nd$-qubit depolarized GHZ state to be advantageous over the optimal local strategy, which also scales as $1/d$ and is almost independent of $n$. Imperfect local entanglement generation will decrease the GHZ fidelity to the threshold at $n_{\mathrm{max}}$. Therefore, we expect $k^{n_{\mathrm{max}}-1}F\approx 1/d$, which gives $n_{\mathrm{max}}\approx-\ln(dF)/\ln(k)$. This also implies robustness of the quantum advantage in DQS against the imperfection of probe state generation over realistic quantum networks. See more details in Sec. III.B of the Supplemental Material [54].
Local measurement constraint.—It is possible that the optimal measurement to saturate the QCRB is entangling between the sensor nodes. In the DQS setup, entangling measurement needs additional remote entanglement as a resource to implement. Therefore, we want to use only local operations and classical communication (LOCC) [68] to extract information in DQS problems. However, in general LOCC is not guaranteed to saturate the QCRB in DQS [69]. Here we explicitly consider the local measurement constraint.
It is known that $M=\sigma_{x}^{\otimes nd}$ is the optimal observable for $nd$-qubit pure GHZ probe state under unitary z-direction phase accumulation [65], and in principle each sensor node only needs to perform measurement of the local observable $\sigma_{x}^{\otimes n}$. However, if we use this measurement for noisy GHZ states, for instance depolarized GHZ state, the variance of parameter estimation diverges in the limit of small local parameters. We note that this conclusion is also implied by numerical results in another recent work by Cao and Wu [70]. In Sec. IV.A of the Supplemental Material [54] we analytically demonstrate this property for the depolarized GHZ state, while it holds generally for any GHZ-diagonal state with non-unit fidelity.
Motivated by our problem formulation that the parameter to estimate is encoded through z-direction phase accumulation, we further explore the optimization over a subset of local measurements, i.e. the tensor product of single-qubit measurements along a direction on the equator of the Bloch sphere: $M(\alpha)=\left[O(\alpha)\right]^{\otimes nd}$ where $O(\alpha)=|\psi^{+}(\alpha)\rangle\langle\psi^{+}(\alpha)|-|\psi^{-}(\alpha)\rangle\langle\psi^{-}(\alpha)|$ with $|\psi^{\pm}(\alpha)\rangle=(|0\rangle\pm e^{i\alpha}|1\rangle)$ (thus $O(\alpha)=e^{i\alpha}|1\rangle\langle 0|+e^{-i\alpha}|0\rangle\langle 1|$), characterized by the azimuthal angle $\alpha$. For an $nd$-qubit depolarized GHZ state, the optimal azimuthal angle is $\alpha_{\mathrm{opt}}=\frac{2l+1}{2nd}\pi,\ l\in\mathbb{Z}$. Note that the optimal azimuthal angle depends on number of local quantum sensors $n$ and sensor node number $d$. Moreover, the estimation variance diverges quickly when $\alpha$ deviates from $\alpha_{\mathrm{opt}}$. This implies that the accuracy of local operation is extremely important, and potentially more important than the quality of the entangled states distributed by quantum networks.
Figure 2: Visualization of fidelity thresholds. The blue curve denotes the threshold given by the QFI, and the red curve denotes the threshold given by optimized local azimuthal measurement. The horizontal and curved black dashed lines represent the thresholds for $d$-qubit depolarized GHZ state to be GME and non-fully separable, respectively.
The fidelity threshold for an $nd$-qubit depolarized GHZ state to be advantageous over the optimal local strategy when using the optimized azimuthal measurement is:
$F_{\mathrm{th},M(\alpha_{\mathrm{opt}})}(n)=\frac{2^{nd}+\sqrt{d}-1}{2^{nd}\sqrt{d}}.$ (5)
The case of $n=1$ for Eq. 5 corresponds to noiseless local entanglement generation, similar to Eq. 4. We thus visualize and compare both fidelity thresholds in Fig. 2. The fidelity threshold in Eq. 5 is always higher than that given by Eq. 4, suggesting that the optimized local azimuthal measurement does not saturate the QCRB. On the other hand, although for small problem sizes $d<5$ the initial GHZ state needs to be GME to demonstrate quantum advantage if we use the optimized azimuthal measurement, when the problem size grows the requirement of initial fidelity drops, and in general the initial $d$-qubit GHZ state does not have to be GME. On the other hand, we can also show that the optimized local azimuthal measurement can saturate the QCRB for rank-2 dephased GHZ states, which implies an error model-dependent separability of optimal measurement for DQS. More details on the azimuthal measurement optimization can be found in Sec. IV.B of the Supplemental Material [54]. Moreover, we can use the GHZ state fidelity thresholds to estimate Bell state fidelity requirement, see Sec. IV.C of the Supplemental Material [54].
Table 1: Simulation results for the three scenarios.
$p$ $\eta$ $\tilde{\eta}$ $F$
Scenario 1 $\approx 0.02$ $\approx 0.95<1$ $\approx 0.02<1$ 0.591
Scenario 2 $\approx 0.72$ $\approx 1.57>1$ $\approx 1.13>1$ 0.732
Scenario 3 $\approx 0.81$ $\approx 2.19>1$ $\approx 1.77>1$ 0.854
Quantum network simulation.—Following the above analytics, we simulate GHZ state distribution given realistic first-generation quantum repeater [71, 72, 73, 74] protocol stacks with an open-source, customizable quantum network simulator, SeQUeNCe [75]. We consider a simple 3-node network with linear topology, where a center node directly connects the other two end nodes through optical fibers, with a Bell state measurement station in the middle of each fiber link. First, bipartite entanglement links are established between the center node and the other nodes. Then LOCC, such as gate teleportation [76, 77, 78] and graph state fusion [79, 80], are performed to generate a GHZ state across all network nodes.
In our simulations we consider three scenarios, and thus three sets of system parameter values: $\{0.01\mathrm{s},0.1\mathrm{s},1\mathrm{s}\}$ for memory coherence times, $\{0.05,0.1,0.5\}$ for memory efficiency, and $\{0.8,0.85,0.9\}$ for raw entanglement fidelity. The first simulation scenario uses the first value in each of the above sets, and so on. Meanwhile, we fix other system parameters, especially: a 1s time interval allowed for entanglement distribution, a 10km distance between the center node and end nodes, 10 memories per end node and 20 memories on the center node so that entanglement purification [81, 82] is possible. We emphasize that the parameter values are not chosen from a specific reported experiment, but are selected according to the general vision of the state of the art and the potential future development of various candidate platforms for quantum networking, including solid state systems [39, 41, 83], atomic systems [40, 84], and superconducting systems [85]. The network topology is visualized in Fig. 3. Details of our quantum network simulation can be found in Sec. VII of the Supplemental Material [54].
Figure 3: Topology of the simulated 3-node quantum network.
We characterize the performance of probe state distribution by three figures of merit, namely the success probability of distributing the 3-qubit GHZ state, $p$, the relative advantage, $\eta$, and the normalized relative advantage, $\tilde{\eta}=p\cdot\eta$, which takes into account failure in entanglement distribution (See Sec. V of the Supplemental Material [54]). For each scenario we repeat the simulation 1000 times, and $\eta$ is calculated based on the average of density matrices of successfully distributed GHZ states under ensemble interpretation. The results are collected in Table 1, from which it is clear that the most modest parameter choice does not permit quantum advantage, while when hardware performance improves quantum advantage becomes possible without changing the realistic quantum network protocol stack. We also note the average fidelity $F$ of the successfully generated states for each scenario; as expected, the fidelity increases as the network parameters improve.
Conclusion and discussion.—In summary, we perform extensive analytical studies on the impact of imperfect GHZ state distributed by noisy quantum networks on DQS. Our results offer new insights into realistic DQS, and reveal the relation between entanglement and DQS quantum advantage. We also simulate the GHZ state distribution process over a 3-node quantum network, demonstrating the possibility of DQS quantum advantage with realistic quantum network stacks. The new features we develop in SeQUeNCe to reflect imperfections in entanglement distribution are completely open-source [86], and can thus serve as valuable resources for future quantum network research. We leave more detailed quantum network simulation as a followup work.
We have primarily focused on probe state preparation errors because it is the major error source in quantum networks. It could also be interesting to evaluate the impact of sensor decoherence after state preparation [87], where the time dependence of estimation accuracy gain [88, 89, 90] becomes important. We note that quantum error correction [91, 92, 93, 94] offers a good opportunity in fighting against noise in sensing process. Meanwhile, protocols such as continuous entanglement distribution [95, 96, 97, 98, 99] might be necessary to reduce latency of probe state preparation over quantum networks. In addition, optimization of Bell-state-based graph state distribution [100, 101, 102, 103, 104, 105, 106, 107, 108] is important for larger scale DQS. Notably, privacy and security [109, 110, 111] is a potentially important aspect of realistic quantum sensor network as well. Finally, although we mainly assume finite-dimensional matter-based quantum sensors, photonic systems [112, 113, 114, 115, 116, 117] are also playing an important role in DQS.
Acknowledgments.—A.Z. would like to thank Tian-Xing Zheng and Boxuan Zhou for helpful discussions. This material is based upon work supported by the U.S. Department of Energy Office of Science National Quantum Information Science Research Centers. Work performed at the Center for Nanoscale Materials, a U.S. Department of Energy Office of Science User Facility, was supported by the U.S. DOE, Office of Basic Energy Sciences, under Contract No. DE-AC02-06CH11357. A.Z. and T.Z. are also supported by the NSF Quantum Leap Challenge Institute for Hybrid Quantum Architectures and Networks (NSF Grant No. 2016136), and the Marshall and Arlene Bennett Family Research Program.
Supplemental Material
I Multiparameter estimation
In this section we provide the background information for distributed quantum sensing problem, including multiparameter estimation, classical and quantum Fisher information matrix, classical and quantum Cramér-Rao bound, and estimating linear function of local parameters.
I.1 (Quantum) Cramér-Rao bound
In a general multiparameter estimation problem, there are $d$ parameters $x=(x_{1},\dots,x_{d})^{T}\in\mathbb{R}^{d}$ to be estimated. To perform the estimation, a collection of statistical data (sample) $s=(s_{1},\dots,s_{\mu})^{T}\in\mathbb{R}^{\mu}$ are accumulated from $\mu$ times of measurement. In the so-called probabilistic approach of estimation, it is assumed that the measured data are random numbers subject to a probability distribution $p(s|x)$ which is dependent on the parameters. An estimator $\hat{x}$ is a rule for calculating the estimate value of parameters based on the measured data. For an estimator, its quality can be characterized by the covariance matrix:
$\left(\mathrm{Cov}(\hat{x})\right)_{ij}=\mathbb{E}\left[(\hat{x}_{i}-\mathbb{E}[\hat{x}_{i}])(\hat{x}_{j}-\mathbb{E}[\hat{x}_{j}])\right],$ (6)
where $\mathbb{E}[\cdot]$ denotes the estimation value of a certain random variable. It is commonly assumed that the estimator is locally unbiased, i.e. $\mathbb{E}[\hat{x}_{i}]=x_{i}$. It is well known that the covariance matrix of any locally unbiased estimator obeys the Cramér-Rao bound (CRB):
$\mathrm{Cov}(\hat{x})\geq\frac{\mathcal{I}^{-1}}{\mu},$ (7)
where for matrices $A\geq B$ means that the matrix difference $(A-B)$ is positive semidefinite, $N$ is the number of data points in the sample, and $\mathcal{I}$ is the classical Fisher information matrix (CFIM):
$\mathcal{I}_{ij}=\mathbb{E}\left[\frac{\partial\ln p(s|x)}{\partial x_{i}}\frac{\partial\ln p(s|x)}{\partial x_{j}}\right].$ (8)
Note that the CRB holds when the CFIM is invertible, i.e. positive definite given its symmetric nature.
When performing multiparameter estimation using quantum mechanical systems, the parameters are usually encoded in the quantum state $\rho_{x}$. The process of obtaining data points through measurement can then be described by POVMs $\{\Pi_{s}\}$ which satisfy normalization $\sum_{s}(\int_{s})\Pi_{s}=I$. That is, now the probability distribution of sample data points is $p(s|x)=\mathrm{tr}(\rho_{x}\Pi_{s})$, from which the CFIM can be calculated.
The quantum Fisher information matrix (QFIM) $\mathcal{F}$ is defined through symmetric logarithmic derivatives (SLD) with respect to different single parameters:
$\mathcal{F}_{ij}=\mathrm{Tr}\left(\rho\frac{L_{i}L_{j}+L_{j}L_{i}}{2}\right),$ (9)
where $L_{i}$ is the SLD for parameter $x_{i}$, s.t.
$\frac{\partial}{\partial x_{i}}\rho\coloneqq\partial_{i}\rho=\frac{1}{2}\{\rho,L_{i}\}.$ (10)
The QFIM satisfies $\mathcal{F}\geq\mathcal{I}_{\Pi}$, for arbitrary POVM $\{\Pi\}$ applied on $\rho_{x}$. Therefore, we can use the QFIM to bound the covariance matrix of any locally unbiased estimator $\hat{x}$:
$\mathrm{Cov}(\hat{x})\geq\frac{\mathcal{I}^{-1}}{\mu}\geq\frac{\mathcal{F}^{-1}}{\mu},$ (11)
which is known as the quantum Cramér-Rao bound (QCRB).
I.2 Estimating functions of parameters
Besides “naturally” encoded parameters, we might want to estimate functions of them. That is, we can consider $d$ new parameters $\theta=(f_{1}(x),\dots,f_{d}(x))^{T}\in\mathbb{R}^{d}$. It turns out that the QFIM of the derived parameters $\theta$ can be expressed in terms of the QFIM of the original parameters $x$ as [61]:
$\mathcal{F}(\theta)=J^{T}\mathcal{F}(x)J,$ (12)
where $J$ is the Jacobian matrix whose matrix elements are $J_{ij}=\partial x_{i}/\partial\theta_{j}$. Then the QCRB for the estimator $\hat{\theta}$ of the derived parameters becomes:
$\mathrm{Cov}(\hat{\theta})\geq\frac{\mathcal{F}^{-1}(\theta)}{\mu}.$ (13)
In some cases such as this work, we may focus on linear functions of parameters:
$\theta=Mx,$ (14)
where without loss of generality we require that all rows of $M$ are linearly independent. Then the Jacobian is simply $J=M^{-1}$. Now consider the case where we are only interested in one specific parameter as a linear function of $d$ parameters, $\theta_{1}=v_{1}^{T}x$. We may construct $(d-1)$ additional vectors $v_{i}$, $i=2,\dots,d$ that are mutually linearly independent, and also linearly independent of $v_{1}$. Then we construct $M=(v_{1},\dots,v_{d})^{T}$ which determines $n$ derived parameters, with $\theta_{1}=(Mx)_{1}$. The variance of estimating the single derived parameter $\theta_{1}$ is then bounded from the QCRB [61]:
$\mathrm{Var}(\hat{\theta}_{1})\geq\frac{\left(\mathcal{F}^{-1}(\theta)\right)_{11}}{\mu}=\frac{\left(J^{-1}\mathcal{F}^{-1}(x)(J^{T})^{-1}\right)_{11}}{\mu},$ (15)
at fixed values of other derived parameters $\theta_{2},\dots,\theta_{n}$, which is valid in that the additional parameters are constructed to be linearly independent of the single parameter of our interest.
II Quantum Fisher information matrix under unitary encoding
Suppose the encoded state has spectral decomposition $\rho=\sum_{a}\lambda_{a}|\psi_{a}\rangle\langle\psi_{a}|$ with all $\lambda_{a}$ being non-zero. Then the QFIM can be expressed as [64]
$\mathcal{F}_{ij}=$ $\sum_{a}\frac{(\partial_{i}\lambda_{a})(\partial_{j}\lambda_{a})}{\lambda_{a}}+\sum_{a}4\lambda_{a}\mathrm{Re}(\langle\partial_{i}\psi_{a}|\partial_{j}\psi_{a}\rangle)-\sum_{a,b}\frac{8\lambda_{a}\lambda_{b}}{\lambda_{a}+\lambda_{b}}\mathrm{Re}(\langle\partial_{i}\psi_{a}|\psi_{b}\rangle\langle\psi_{b}|\partial_{j}\psi_{a}\rangle).$ (16)
We note that there are different equivalent ways of calculating QFIM, as documented in some other extensive reviews, for instance [61, 62, 63].
Then we consider that the parameter dependence of $\rho$ comes from unitary encoding $U(x)$ of an $x$-independent initial state $\rho_{0}=\sum_{a}\lambda_{a0}|\psi_{a0}\rangle\langle\psi_{a0}|$, i.e. $\rho=U(x)\rho_{0}U(x)^{\dagger}$. Now the QFIM can be re-expressed in the convention of [64] as:
$\mathcal{F}_{ij}=\sum_{a}4\lambda_{a0}\mathrm{cov}_{|\psi_{a0}\rangle}(G_{i},G_{j})-\sum_{a\neq b}\frac{8\lambda_{a0}\lambda_{b0}}{\lambda_{a0}+\lambda_{b0}}\mathrm{Re}(\langle\psi_{a0}|G_{i}|\psi_{b0}\rangle\langle\psi_{b0}|G_{j}|\psi_{a0}\rangle),$ (17)
where $\mathrm{cov}_{|\psi\rangle}(A,B)$ denotes the covariance of observables $A$ and $B$ under a pure state $|\psi\rangle$:
$\mathrm{cov}_{|\psi\rangle}(A,B)=\frac{1}{2}\langle\psi|\{A,B\}|\psi\rangle-\langle\psi|A|\psi\rangle\langle\psi|B|\psi\rangle,$ (18)
and $G_{i}$ is the generator of parameter $x_{i}$:
$G_{i}=i(\partial_{i}U^{\dagger})U=-iU^{\dagger}(\partial_{i}U).$ (19)
II.1 GHZ-diagonal state with collective spin phase accumulation
II.1.1 Justification for considering GHZ-diagonal initial state
At the beginning of this section, we elaborate on the mechanisms that can in principle allow us to have GHZ-diagonal states as the initial probe for DQS.
Pauli twirling.— Pauli twirling [55, 56, 57] can be utilized in two steps during the GHZ state distribution by the quantum network to make the final state in a GHZ-diagonal form. First of all, we distribute 2-qubit Bell states between quantum sensor nodes, which could in principle be non-Bell-diagonal. However, for each Bell state we can perform bilocal Pauli twirling: $\rho\to\sum_{i=1}^{4}(P_{i}\otimes P_{i})\rho(P_{i}\otimes P_{i})/4$ where $P_{i}$ are standard 1-qubit Pauli operators and the summation is over all 4 Pauli operators. It is easy to check that the resulting state is in Bell-diagonal form, i.e. pure Bell state undergoing Pauli channel. Then, we need to assemble the Bell states together to create the global GHZ state (as reviewed in Sec. VI). The assembly process involves local Clifford operations especially entangling gates which are inevitably noisy in practice. Remarkably, the noise introduced by quantum gates, and in general quantum circuits that can be compiled into layers of quantum gates, can be effectively transformed into incoherent errors with Pauli channel an example, using techniques such as randomized compiling [118, 119]. As a result, since the assembly process only applies Clifford operations and introduces Pauli errors to Bell-diagonal states, the final state is effectively a pure GHZ state undergoing a Pauli channel, which is in GHZ-diagonal form. The twirling operations are also all local, so is in principle applicable to quantum network scenarios.
GHZ stabilizer twirling.— On the other hand, we can use GHZ stabilizer twirling [58] at the very end of GHZ state distribution, to turn any distributed global GHZ state into GHZ-diagonal form without changing the diagonal elements in the density matrix. Explicitly, the GHZ state stabilizer twirling is an process of randomly applying the elements of the stabilizer group $\mathcal{G}=\{g_{i}\}$ for the GHZ state $\rho\to\sum_{i}g_{i}\rho g_{i}/|\mathcal{G}|$, where we have considered that stabilizer group elements are all Pauli strings so we have dropped the Hermitian conjugate. For completeness, we give a simple proof of why such GHZ stabilizer twirling is able to depolarize any state into GHZ-diagonal form without changing the diagonal elements.
Proposition II.1.
GHZ stabilizer twirling turns any state into GHZ-diagonal form with diagonal elements unchanged.
Proof.
For an arbitrary $n$-qubit state $\rho$, we can always represent its density matrix in the GHZ basis. Now we explicitly apply the GHZ stabilizer twirling on $\rho$.
(1) Diagonal elements of the density matrix in GHZ basis. For the element corresponding to the standard GHZ state this follows directly from the definition of stabilizer. Other GHZ basis states $|\psi\rangle$ can be obtained through applying Pauli strings to the standard GHZ state $|\psi\rangle=P|\mathrm{GHZ}\rangle$, where $P$ is a certain $n$-qubit Pauli string. Then the application of GHZ stabilizer elements on $|\psi\rangle$ gives
$g_{i}|\psi\rangle=g_{i}P|\mathrm{GHZ}\rangle=\left(Pg_{i}-[P,g_{i}]\right)|\mathrm{GHZ}\rangle=(-1)^{f([P,g_{i}])}P|\mathrm{GHZ}\rangle=(-1)^{f([P,g_{i}])}|\psi\rangle,$ (20)
where we define $f(\cdot)$ to output zero when the input is zero and output one otherwise, and we have used the fact that Pauli strings either commute or anti-commute. In other words, the application of GHZ stabilizers on a GHZ basis state gives either the basis state itself or with an additional $(-1)$ factor. For GHZ basis states $|\psi\rangle$ we have $g_{i}|\psi\rangle\langle\psi|g_{i}=\left[(-1)^{f([P,g_{i}])}\right]^{2}|\psi\rangle\langle\psi|=|\psi\rangle\langle\psi|$. Therefore, any diagonal element of the density matrix in GHZ basis will not be changed under GHZ stabilizer twirling.
(2) off-diagonal elements. Recall the the stabilizer group is Abelian and the $n$-qubit GHZ stabilizer group is generated by $n$ commuting $n$-qubit Pauli strings. Therefore, the additional factor, i.e. $(+1)$ or $(-1)$, coming from applying GHZ stabilizer element to GHZ basis states can be determined by the additional factor obtained from applying the GHZ stabilizer generators. For each GHZ basis state $|\psi_{j}\rangle$ where $j=1,\dots,2^{n}$ is the index, we denote the set of generators which give plus one factor $\mathcal{P}_{j}$ and the set of generators which give minus one factor $\mathcal{M}_{j}$, such that $|\mathcal{P}_{j}|+|\mathcal{M}_{j}|=n$. Note that for different $j$ it is impossible to have identical $\mathcal{P}_{j}$ and $\mathcal{M}_{j}$, because if so the two states are identical and should have the same index.
We have that the application of half of the stabilizer elements will result in additional $(-1)$ factor, while the other half will result in $(+1)$ factor. This can be seen by considering $\mathcal{P}_{j}$ and $\mathcal{M}_{j}$, and let $|\mathcal{P}_{j}|=p_{j}$ and $|\mathcal{M}_{j}|=m_{j}$. The number of stabilizer elements that result in additional $(-1)$ factor can be calculated as
$\left.2^{p_{j}}\sum_{i=0}^{\frac{m-1}{2}}\binom{m_{j}}{2i+1}\right|_{\mathrm{odd}~{}m_{j}}=\left.2^{p_{j}}\sum_{i=0}^{\frac{m-2}{2}}\binom{m_{j}}{2i+1}\right|_{\mathrm{even}~{}m_{j}}=2^{m_{j}+p_{j}-1}=2^{n-1}=\frac{|\mathcal{G}|}{2}.$ (21)
This directly implies that $|\psi_{j}\rangle\langle\mathrm{GHZ}|$ and $|\mathrm{GHZ}\rangle\langle\psi_{j}|$ will vanish after GHZ stabilizer twirling due to the cancellation of $(-1)$ and $(+1)$ factor terms. For the application of GHZ stabilizer twirling to general $|\psi_{j}\rangle\langle\psi_{k}|$, we consider that arbitrary two GHZ basis states can be transformed into each other via application of a Pauli string. Therefore, we have
$g_{i}|\psi_{j}\rangle\langle\psi_{k}|g_{i}=g_{i}|\psi_{j}\rangle\langle\psi_{j}|Pg_{i}=(-1)^{f([P,g_{i}])}g_{i}|\psi_{j}\rangle\langle\psi_{j}|g_{i}P=(-1)^{f([P,g_{i}])}|\psi_{j}\rangle\langle\psi_{j}|P=(-1)^{f([P,g_{i}])}|\psi_{j}\rangle\langle\psi_{k}|.$ (22)
Consequently, the above argument of evaluating the additional factor introduced from the application of GHZ stabilizer element still applies. We thus know that again half of the stabilizer elements give $f([P,g_{i}])=0$ while the other half gives $f([P,g_{i}])=1$. In the end, under GHZ stabilizer twirling any $|\psi_{j}\rangle\langle\psi_{k}|$ vanishes as long as $j\neq k$. ∎
The proof applies to arbitrary graph states, and thus any state can be twirled to be diagonal in arbitrary graph state basis. Moreover, the stabilizer elements are simply Pauli strings, so the stabilizer twirling can in principle be implemented over quantum networks.
II.1.2 Derivation of quantum Fisher information matrix
Consider that the parameter encoding is through $U(x)=\exp\left[-i\left(\sum_{i=1}^{d}x_{i}H_{i}\right)\right]$, where $d$ denotes the total number of sensor nodes in the network, $H_{i}=\frac{1}{2}\sum_{k=0}^{n-1}\sigma^{(i,k)}_{z}$ is the local collective spin of $n$ qubit sensors on node $i$, and the parameters $x_{i}$ physically correspond to accumulated phases through Hamiltonian evolution. It can thus be easily verified that the generators according to the above definition are $G_{i}=-H_{i}$, where the negative sign does not matter as generators appear in pairs in Eq. 17. Then for $\rho_{0}$ which is a GHZ-diagonal state, we have that:
$\mathrm{cov}_{|\psi_{m0}\rangle}(G_{i},G_{j})=$ $\mathrm{cov}_{|\psi_{m0}\rangle}(H_{i},H_{j})=\frac{1}{2}\langle\psi_{m0}|\{H_{i},H_{j}\}|\psi_{m0}\rangle,\ \forall m,$ (23)
because weight-1 Pauli strings do not stabilize the standard GHZ states. Meanwhile, we have that:
$\langle\psi|\sigma_{z}^{(a)}\sigma_{z}^{(b)}|\psi\rangle=1,\ \forall a,b,$ (24)
where $a,b$ are indices for the qubits in $|\psi\rangle$ which is a GHZ-basis state. This is because weight-2 Pauli strings with only Pauli Z operators stabilize the standard GHZ states. Therefore, for any GHZ-diagonal $\rho_{0}$, the first single-index sum in Eq. 17 is always a constant:
$\sum_{a}4\lambda_{a0}\mathrm{cov}_{|\psi_{a0}\rangle}(G_{i},G_{j})=$ $\sum_{a}4\lambda_{a0}\mathrm{cov}_{|\psi_{a0}\rangle}\left(\frac{1}{2}\sum_{k=0}^{n-1}\sigma^{(i,k)}_{z},\frac{1}{2}\sum_{l=0}^{n-1}\sigma^{(j,l)}_{z}\right)$
$=$ $\sum_{a}\lambda_{a0}\sum_{k=0}^{n-1}\sum_{l=0}^{n-1}\mathrm{cov}_{|\psi_{a0}\rangle}\left(\sigma^{(i,k)}_{z},\sigma^{(j,l)}_{z}\right)=n^{2},\ \forall i,j,$ (25)
which reveals the Heisenberg scaling with number of local qubit sensors. Then the derivation of the QFIM reduces to the second term in Eq. 17.
There are $2^{nd}$ orthonormal $nd$-qubit GHZ states across $d$ sensor nodes where each node holds $n$ qubits, which can be labeled by the binary string from 0 through $12^{nd-1}-$ 1: These states are superpositions of two computational basis states which correspond to binary strings of $b\in\{0,1,\dots,2^{nd-1}-1\}$ and $12^{nd}-b-$ 1, respectively. The additional $2^{nd-1}$ states come from adding an additional $\pi$ relative phase between the two computational basis states. Notice that for any GHZ state $|\psi\rangle$, the application of a single Pauli Z operator on it will lead to a $\pi$ change in the relative phase between the two computational basis states in superposition. Therefore, $\langle\psi_{a0}|G_{i}|\psi_{b0}\rangle$ is either zero due to orthogonality between $|\psi_{a0}\rangle$ and $G_{i}|\psi_{b0}\rangle$, or one, and it takes unit value only when $|\psi_{a0}\rangle$ and $|\psi_{b0}\rangle$ correspond to the same length-$nd$ binary string and have relative phases which differ by $\pi$.
For a general $nd$-qubit GHZ-diagonal state $\rho_{0}=\sum_{a}\lambda_{a0}|\psi_{a0}\rangle\langle\psi_{a0}|$ where each node has $n$ qubits, we use $\mathcal{S}$ to denote the set of index pairs $(a,b)$ such that $|\psi_{a0}\rangle$ and $|\psi_{b0}\rangle$ are GHZ states as superposition of the same pair of computational basis states but with opposite relative phase. Note that for such pairs, $(a,b)$ and $(b,a)$ are both included in set $\mathcal{S}$. Then we have:
$\sum_{a\neq b}\frac{8\lambda_{a0}\lambda_{b0}}{\lambda_{a0}+\lambda_{b0}}\mathrm{Re}(\langle\psi_{a0}|G_{i}|\psi_{b0}\rangle\langle\psi_{b0}|G_{j}|\psi_{a0}\rangle)=n^{2}\sum_{(a,b)\in\mathcal{S}}\frac{2\lambda_{a0}\lambda_{b0}}{\lambda_{a0}+\lambda_{b0}}=Cn^{2},\ \forall i,j,$ (26)
which again does not depend on $i,j$, and will only modify the Heisenberg scaling factor, while we comment that in practice $C$ can be dependent on $n$. We may consider a special case as example: We assume noiseless local entanglement generation when extending any $d$-qubit GHZ state into $nd$-qubit GHZ state involving $n-1$ additional qubits per node, and the result is simply duplicating each binary digit of the binary strings that represent the computational basis. For instance, $(|000\rangle+|111\rangle)/\sqrt{2}$ will be extended to $(|\tilde{0}_{n}\tilde{0}_{n}\tilde{0}_{n}\rangle+|\tilde{1}_{n}\tilde{1}_{n}\tilde{1}_{n}\rangle)/\sqrt{2}$, where $\tilde{i}_{n}=\underbrace{i\dots i}_{\text{$n$ digits}}$. In this example $C$ is indeed a constant that only depends on the initial $d$-qubit GHZ state.
We also comment that the form of the series in Eq. 26 implies that only Pauli Z errors affect the QFIM, which is a result of our assumed encoding channel, i.e. a z-axis coupling. Additionally, we note that the maximum value of $C$ is 1 for GHZ-diagonal states, and it is achieved if and only if every pair of states in set $\mathcal{S}$ have identical eigenvalues. This can be seen as follows:
$C=$ $\sum_{(a,b)\in\mathcal{S}}\frac{2\lambda_{a0}\lambda_{b0}}{\lambda_{a0}+\lambda_{b0}}=\frac{1}{2}\sum_{(a,b)\in\mathcal{S}}\frac{(\lambda_{a0}+\lambda_{b0})^{2}-(\lambda_{a0}-\lambda_{b0})^{2}}{\lambda_{a0}+\lambda_{b0}}$
$=$ $\frac{1}{2}\sum_{(a,b)\in\mathcal{S}}(\lambda_{a0}+\lambda_{b0})-\frac{1}{2}\sum_{(a,b)\in\mathcal{S}}\frac{(\lambda_{a0}-\lambda_{b0})^{2}}{\lambda_{a0}+\lambda_{b0}}$
$=$ $1-\frac{1}{2}\sum_{(a,b)\in\mathcal{S}}\frac{(\lambda_{a0}-\lambda_{b0})^{2}}{\lambda_{a0}+\lambda_{b0}},$ (27)
where the subtracted term is non-negative, and it becomes zero if and only if $\lambda_{a0}=\lambda_{b0}$, $\forall(a,b)\in\mathcal{S}$.
Combining the above results, we can explicitly write the QFIM with respect to local parameters $x=(x_{1},\dots,x_{d})^{T}$ as:
$\mathcal{F}(x)=(1-C)n^{2}\begin{pmatrix}1&1&\dots&1\\ 1&1&\dots&1\\ \vdots&\vdots&\ddots&\vdots\\ 1&1&\dots&1\\ \end{pmatrix}.$ (28)
According to the parameter estimation problem of our interest, we have that $v_{1}\propto(1,\dots,1)^{T}\in\mathbb{R}^{d}$. For concreteness, we may choose $v_{1}=(1,\dots,1)^{T}/\sqrt{d}$, which is a normalized vector under 2-norm. As now we only focus on estimating one parameter $v_{1}^{T}x$, we can construct an orthonormal matrix $M=(v_{1},\dots,v_{d})^{T}$ such that $v_{i}^{T}v_{j}=\delta_{ij}$. Then according to Eq. 12 we have the QFIM with respect to new parameters $\theta=(v_{1}^{T}x,\dots,v_{d}^{T}x)$:
$\mathcal{F}(\theta)=$ $\sqrt{d}(1-C)n^{2}\begin{pmatrix}v_{1}^{T}\\ v_{2}^{T}\\ \vdots\\ v_{d}^{T}\\ \end{pmatrix}\begin{pmatrix}v_{1}&v_{1}&\dots&v_{1}\end{pmatrix}\begin{pmatrix}v_{1}&v_{2}&\dots&v_{d}\end{pmatrix}$
$=$ $\sqrt{d}(1-C)n^{2}\begin{pmatrix}1&1&\dots&1\\ 0&0&\dots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\dots&0\\ \end{pmatrix}\begin{pmatrix}v_{1}&v_{2}&\dots&v_{d}\end{pmatrix}$
$=$ $d(1-C)n^{2}\begin{pmatrix}v_{1}^{T}\\ 0\\ \vdots\\ 0\\ \end{pmatrix}\begin{pmatrix}v_{1}&v_{2}&\dots&v_{d}\end{pmatrix}=d(1-C)n^{2}\begin{pmatrix}1&0&\dots&0\\ 0&0&\dots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\dots&0\\ \end{pmatrix}.$ (29)
II.2 Lower bound of QFI for a fixed fidelity
Given the analytical formula of calculating the $C$ term in QFI expression, we could evaluate the lower bound of QFI for all possible GHZ-diagonal states with a fixed fidelity. We could prove the following result.
Proposition II.2.
For $d$-qubit GHZ-diagonal states with a fixed fidelity $F\in(0,1)$, the lowest QFI for estimating the average of $d$ local parameters is $d(2F-1)^{2}$.
Proof.
We label the the eigenvalues of the density matrix in the following way. Consider that GHZ states can be expressed as a superposition of two computational basis states, which correspond to two binary strings, and one of the binary string corresponds to a smaller integer $n\in\{0,1,\dots,2^{d-1}-1\}$ while the other corresponds to a larger integer $m=2^{d}-n-1$. For a GHZ state expressed as a superposition of computational basis state corresponding to $n$ and $m$ with $n<m$, we let its index be $2n$ if the relative phase between two computational basis states is 00, and $12n+$ 1 otherwise. For instance, the standard GHZ state $|\mathrm{GHZ}_{d}\rangle$ has index 0, and $Z|\mathrm{GHZ}_{d}\rangle$ has index 1. Then we have that the eigenvalue corresponding to $|\mathrm{GHZ}_{d}\rangle$ is $\lambda_{00}=F$. For the rest $12^{d}-$ 1 eigenvalues with $i\in\{1,2,\dots,2^{d}-1\}$, we can express them as $\lambda_{i0}=(1-F)p_{i}$, s.t. $p_{i}\in[0,1]$ and $\sum_{i=1}^{2^{d}-1}p_{i}=1$.
Then the objective is transformed to finding the combination of $(p_{1},\dots,p_{2^{d}-1})$ which gives the highest $C$ under the above constraints. The constraints clearly define a closed and compact region $\mathcal{R}$. Then according to the extreme value theorem, we are sure that there exists a maximum value and a minimum value for any continuous function of $(p_{1},\dots,p_{2^{d}-1})$ on $\mathcal{R}$, and the extreme values must be taken either on the boundary of $\mathcal{R}$, or at critical points inside $\mathcal{R}$.
Now $C$ can be re-written as:
$C=\frac{4F(1-F)p_{1}}{F+(1-F)p_{1}}+\sum_{i=1}^{2^{d-1}-1}\frac{4(1-F)p_{2i}p_{2i+1}}{p_{2i}+p_{2i+1}},$ (30)
which is obviously continuous on $\mathcal{R}$. We first take the partial derivatives with respect to $p_{i}$:
$\frac{\partial}{\partial p_{1}}C=\frac{4(1-F)F^{2}}{\left[F+(1-F)p_{1}\right]^{2}}>0,$ (31)
[…]
which means that there is no critical point inside $\mathcal{R}$, so the maximum value can only be on the boundary of $\mathcal{R}$.
It is clear that the boundary of $\mathcal{R}$ can be divided into $12^{d}-$ 1 parts $\{B_{i}\}$, each determined by $p_{i}=0$ for a specific $i\in\{1,2,\dots,2^{d}-1\}$. Under the equality constraint $\sum_{i=1}^{2^{d}-1}p_{i}=1$, the above partial derivatives suggest that there is still no critical point inside any $B_{i}$. This means that the extreme values on the boundary should be on the boundary of $B_{i}$, i.e, $\{B_{ij}\}$, where the subscript means the $j$-th part of the boundary of $B_{i}$. The conclusion of no inside critical point holds until we have reduced the boundary into zero-dimension points, characterized by $p_{i}=1$.
Finally, we can compare the $C$ values for all the $12^{d}-$ 1 choices of $(p_{1},\dots,p_{2^{d}-1})$. It is obvious that if $p_{i}=1$ and $i>1$, $C=0$, and if $p_{1}=1$, $C=4F(1-F)>0$. Therefore, the maximal value of $C$ on $\mathcal{R}$ is $4F(1-F)$ corresponding to $\rho=F|\mathrm{GHZ}_{d}\rangle\langle\mathrm{GHZ}_{d}|+(1-F)Z|\mathrm{GHZ}_{d}\rangle\langle\mathrm{GHZ}_{d}|Z$, which gives the lowest QFI $\mathcal{F}=d(1-C)=d(2F-1)^{2}$. ∎
The physical interpretation of this result is clear. The worst-case scenario can be interpreted as that all probe state preparation error is indistinguishable from the signal of the parameter to estimate, because of our assumption that the parameter is encoded through z-component phase accumulation. On the other hand, the QFI for noisy GHZ state can be equal to that for noiseless GHZ state, and this can be interpreted as that when probe state error can be distinguished from the signal, it is still possible to extract the information encoded in the “noisy” components of the probe state.
II.3 Quantum advantage condition as entanglement criterion
For entanglement detection purpose, our objective is to show that any fully separable initial probe state cannot result in better DQS performance than the optimal local sensing strategy.
Proposition II.3.
Suppose a $d$-qubit state is subject to a unitary parameter encoding of $d$ parameters $x=(x_{1},\dots,x_{d})^{T}$, $U(x)=\exp\left[-i\left(\sum_{i=1}^{d}x_{i}\frac{\sigma^{(i)}_{z}}{2}\right)\right]$. Then the estimation variance of $\theta_{1}=v_{1}^{T}x$ for $v_{1}=(1,\dots,1)^{T}/\sqrt{d}$ using fully separable $d$-qubit state is always no smaller than $1/\mu$, where $\mu$ is the amount of state copies.
Proof.
Firstly, for pure state $|\varphi\rangle$ undergoing unitary parameter encoding whose all generators commute with each other as in our case, the QFIM can be calculated as
$\mathcal{F}_{ij}=4\left[\frac{1}{2}\langle\varphi|\{H_{i},H_{j}\}|\varphi\rangle-\langle\varphi|H_{i}|\varphi\rangle\langle\varphi|H_{j}|\varphi\rangle\right].$ (33)
Now we consider a fully separable $d$-qubit pure state $|\varphi\rangle_{\mathrm{sep}}=\bigotimes_{i=1}^{d}|\varphi_{i}\rangle$. Its QFIM under the aforementioned unitary parameter encoding channel is thus
$\mathcal{F}_{ij}=$ $\frac{1}{2}\langle\varphi|_{\mathrm{sep}}\{H_{i},H_{j}\}|\varphi\rangle_{\mathrm{sep}}-\langle\varphi|_{\mathrm{sep}}H_{i}|\varphi\rangle_{\mathrm{sep}}\langle\varphi|_{\mathrm{sep}}H_{j}|\varphi\rangle_{\mathrm{sep}}$
$=$ $\left[1-\left(\langle\varphi_{i}|\sigma_{z}|\varphi_{i}\rangle\right)^{2}\right]\delta_{ij}\leq\delta_{ij}.$ (34)
Therefore, we have that for any fully separable $d$-qubit pure state $\mathcal{F}\leq I$, which should be read as $I-\mathcal{F}$ is positive semidefinite.
Then we consider arbitrary fully separable $d$-qubit state that can be expressed as a convex combination of tensor products of subsystem quantum states $\rho_{\mathrm{sep}}=\sum_{i}p_{i}\rho_{i}^{(1)}\otimes\rho_{i}^{(2)}\otimes\dots\otimes\rho_{i}^{(d)}$, where $p_{i}\geq 0,~{}\sum_{i}p_{i}=1$, and in our case $\rho_{i}^{(j)}$ is arbitrary single-qubit density matrix. Notice that every $\rho_{i}^{(j)}$ can be decomposed as a convex combination of single-qubit pure state, so any fully separable state can be decomposed as a convex combination of $d$-qubit fully separable pure states $\rho_{\mathrm{sep}}=\sum_{j}q_{j}|\varphi_{j}^{(1)}\rangle\langle\varphi_{j}^{(1)}|\otimes\dots\otimes|\varphi_{j}^{(d)}\rangle\langle\varphi_{j}^{(d)}|=\sum_{j}q_{j}|\varphi_{j}\rangle\langle\varphi_{j}|_{\mathrm{sep}}$, where $q_{j}\geq 0$ and $\sum_{j}q_{j}=1$.
Recall that the QFIM is convex [64], so we have
$\mathcal{F}(\rho_{\mathrm{sep}})=\mathcal{F}\left(\sum_{j}q_{j}|\varphi_{j}\rangle\langle\varphi_{j}|_{\mathrm{sep}}\right)\leq\sum_{j}q_{j}\mathcal{F}(|\varphi_{j}\rangle\langle\varphi_{j}|_{\mathrm{sep}})\leq I.$ (35)
Also, since QFIM is real symmetric and positive semidefinite, we have the orthogonal decomposition $\mathcal{F}(\rho_{\mathrm{sep}})=Q\Lambda Q^{T}$, where $QQ^{T}=Q^{T}Q=I$ and $\Lambda$ is diagonal s.t. $0\leq\Lambda\leq I$. As before, we want to transform from the current basis of “natural” parameters to another orthonormal basis including $\theta_{1}=v_{1}^{T}x$, through an orthogonal transform $M$ of $\mathcal{F}(\rho_{\mathrm{sep}})$
~$\tilde{\mathcal{F}}(\rho_{\mathrm{sep}})=M\mathcal{F}(\rho_{\mathrm{sep}})M^{T}=MQ\Lambda Q^{T}M^{T}=O\Lambda O^{T},$ (36)
where we use $\tilde{\mathcal{F}}$ to denote the transformed QFIM for the new parameters, and $O=MQ$ is another orthogonal transform s.t. $OO^{T}=O^{T}O=I$.
For estimation of a single parameter in a multiparameter scenario, we have the following bound [1]
$\mathrm{Var}(\hat{\theta}_{1})\geq\frac{\left(\tilde{\mathcal{F}}^{-1}(\rho_{\mathrm{sep}})\right)_{11}}{\mu}\geq\frac{1}{\mu\left(\tilde{\mathcal{F}}(\rho_{\mathrm{sep}})\right)_{11}},$ (37)
where the first inequality is always saturable when we only focus on $\theta_{1}$, while the second, though simpler, is not always saturable. Then we explicitly evaluate (~$\left(\tilde{\mathcal{F}}(\rho_{\mathrm{sep}})\right)_{11}$ to bound it from the above
(~$\left(\tilde{\mathcal{F}}(\rho_{\mathrm{sep}})\right)_{11}=\sum_{i=1}^{d}\lambda_{i}\left(w_{1}^{(i)}\right)^{2}\leq\sum_{i=1}^{d}\left(w_{1}^{(i)}\right)^{2}=w_{1}^{T}w_{1}=1,$ (38)
where we have written $O=(w_{1},\dots,w_{d})^{T}$ with $w_{i}^{T}w_{j}=\delta_{ij}$, $w_{i}^{(i)}$ is the $j$-th element of $w_{i}$, and $\lambda_{i}$ are the diagonal elements of $\Lambda$. From the upper bound of (~$\left(\tilde{\mathcal{F}}(\rho_{\mathrm{sep}})\right)_{11}$ the desired lower bound of $\theta_{1}$ estimation variance is then obvious. ∎
Notice that $1/\mu$ is exactly the estimation variance of the optimal local sensing strategy. The proof can be straightforwardly extended to the scenario where each sensor contains $n$ qubits, and show that any fully separable state with respect to the partition of $d$ sensors cannot achieve $\theta_{1}$ estimation variance lower than $1/(n^{2}\mu)$.
The above thus proves that entanglement in the initial probe state across sensors is necessary to achieve quantum advantage in the DQS task of estimating $\theta_{1}$ over the optimal local sensing strategy. In other words, if we find that a certain input state can demonstrate DQS quantum advantage, there must be some entanglement in it with respect to the partition of $d$ sensors. Although the calculation of $C$ in the quantum advantage condition $d(1-C)>1$ is based on the GHZ-diagonal assumption, as we have demonstrated, arbitrary state can be converted into GHZ-diagonal from through GHZ stabilizer twirling which is local and thus will not generate any entanglement. From this perspective, the quantum advantage condition $d(1-C)>1$ is a sufficient condition for entanglement detection [120, 121] in $d$-partite state where each party has local dimension $2^{n}$, with clear operational interpretation. Importantly, the only calculation needed is $C$, and the calculation is optimization-free. Moreover, even though there are admittedly exponentially many diagonal elements in the multipartite density matrix, in many cases we do not need to go through all of them, because $1-C$ can be rewritten as
$1-C=\sum_{(a,b)\in\mathcal{S}}\frac{(\lambda_{a0}-\lambda_{b0})^{2}}{\lambda_{a0}+\lambda_{b0}},$ (39)
where we have removed double counting in $\mathcal{S}$. Notice that the right hand side is a summation of non-negative terms. Therefore, we can stop the summation when we already have $d\sum_{(a,b)\in\mathcal{S}^{\prime}}\frac{(\lambda_{a0}-\lambda_{b0})^{2}}{\lambda_{a0}+\lambda_{b0}}>1$, where $\mathcal{S}^{\prime}\subset\mathcal{S}$.
II.3.1 An example of non-fully separable state detected by DQS quantum advantage condition
Here we demonstrate one example of $d$-qubit state which is not fully separable and can be detected by the DQS quantum advantage condition. The example state is the equal mixture of GHZ states which can be expressed as superposition of two computational basis states with zero relative phase. For instance, for 3 qubits such GHZ states include $(|000\rangle+|111\rangle)/\sqrt{2},(|001\rangle+|110\rangle)/\sqrt{2},(|010\rangle+|101\rangle)/\sqrt{2},(|011\rangle+|100\rangle)/\sqrt{2}$. For $d$ qubits there are $2^{d-1}$ such GHZ states $|\mathrm{GHZ}_{d}^{(i)}\rangle$ labeled by $i=0,1,\dots,2^{d-1}-1$: The binary string corresponding to one of the two computational basis states equals to the binary representation of the index $i$, while the other computational basis state corresponds to the binary representation of $2^{d}-i$. Then the example state can be written as
$\rho_{d}=\frac{1}{2^{d-1}}\sum_{i=0}^{2^{d-1}-1}|\mathrm{GHZ}_{d}^{(i)}\rangle\langle\mathrm{GHZ}_{d}^{(i)}|.$ (40)
According to the calculation of $(1-C)$ in the DQS quantum advantage condition we have $11-C(\rho_{d})=$ 1, which means that $d[1-C(\rho_{d})]=d>1$, and thus $\rho_{d}$ must be entangled.
The DQS quantum advantage condition easily determines the entanglement in $\rho$, but for other entanglement criteria it could be less straightforward, if possible to successfully detect the entanglement. For instance, according to Laskowski-Żukowski (LZ) criterion [122] for $k$-separability which is equivalent to Mermin-type inequalities [123, 124, 125], a $d$-qubit state that is $d$-separable (fully separable) must satisfy $\max_{j}\left|\rho_{j,\overline{j}}\right|\leq 2^{-d}$, where $\rho_{j,\overline{j}}$ is the off-diagonal element for the density matrix with $j=1,\dots,2^{d}$ being the row index and $\overline{j}=2^{d}-j+1$. This criterion is basis independent, so if a state under a certain basis violates the necessary condition for full separability, it is entangled. However, to obtain the maximum of density matrix off-diagonal element over all possible bases requires additional optimization. It is obvious that the example state $\rho_{d}$ does not violate the $d$-separability necessary condition in two common bases, namely the computational basis and the GHZ basis, so other bases have to be checked for potential detection of entanglement in $\rho_{d}$ with the LZ criterion. In addition, after Dür-Cirac depolarization [66] the example state $\rho_{d}$ becomes separable under arbitrary bipartition, so the Dür-Cirac separability criterion cannot be used to detect entanglement in $\rho_{d}$. Moreover, it can be easily verified for $\rho_{3}$ that it is positive after partial transpose (PPT) under any bipartition, which means that the DQS quantum advantage condition can detect entanglement when the PPT criterion [126, 127] fails.
III Analytics for depolarized GHZ state
III.1 Noiseless local entanglement generation
We consider $d$-qubit GHZ states under collective depolarizing channel which is characterized by a single parameter, the fidelity $F$: $\rho_{d,0}(F)=\frac{2^{d}F-1}{2^{d}-1}|\mathrm{GHZ}_{d}\rangle\langle\mathrm{GHZ}_{d}|+\frac{1-F}{2^{d}-1}I$. For this specific family of states, we can obtain the closed form expression of the constant $C$ in Eq. 26:
$C_{\mathrm{dp}}=$ $2\left[\frac{2F\frac{1-F}{2^{d}-1}}{F+\frac{1-F}{2^{d}-1}}+(2^{d-1}-1)\frac{2\left(\frac{1-F}{2^{d}-1}\right)^{2}}{2\frac{1-F}{2^{d}-1}}\right]=\frac{(1-F)(4^{d}F-2^{d}-2)}{[(2^{d}-2)F+1](2^{d}-1)}.$ (41)
It is obvious that both the numerator and the denominator are positive, and then we take the difference between the numerator and the denominator:
$(1-F)(4^{d}F-2^{d}-2)-[(2^{d}-2)F+1](2^{d}-1)=-(2^{d}F-1)^{2}<0,$ (42)
which means that $10\leq C_{\mathrm{dp}}<$ 1. Then we examine $\eta_{\mathrm{dp}}=d(1-C_{\mathrm{dp}})$:
$\frac{\partial}{\partial F}\eta_{\mathrm{dp}}=\frac{d\left[2-3\times 2^{d}+2^{2d+1}F+(2^{d}-2)4^{d}F^{2}\right]}{(2^{d}-1)\left[(2^{d}-2)F+1\right]^{2}},$ (43)
whose sign is only determined by the numerator. It is easy to find that the above partial derivative equals zero at $F=2^{-d}$ and $F=(2-3\times 2^{d})/(4^{d}-2^{d+1})<0$. That is, the partial derivative is negative for $F\in[0,2^{-d})$ and positive for $F\in(2^{-d},1]$, which means that for fixed $d$ the minimal value of $\eta_{\mathrm{dp}}$ is taken at $F=2^{-d}$. This fidelity corresponds to a $d$-qubit maximally mixed state, and the minimal value is:
$\left.\eta_{\mathrm{dp}}\right|_{F=2^{-d}}=0.$ (44)
This is because unitary encoding does not vary the maximally mixed state and then any following measurement will not be able to extract information of the encoded parameters.
We solve the equation $C_{\mathrm{dp}}=(d-1)/d$, whose only positive solution is the fidelity threshold for depolarized GHZ states to demonstrate advantage in estimating the average of local parameters over the optimal local strategy:
$F_{\mathrm{th,dp}}=2^{-d}+\frac{(2^{d}-1)\left(2^{d}-2+\sqrt{(2^{d}-2)^{2}+2^{d+3}d}\right)}{2^{2d+1}d}.$ (45)
It can be shown that:
Proposition III.1.
$F_{\mathrm{th,dp}}>3/(2^{d}+2)$ for $d\geq 2$.
Proof.
First notice that $3/(2^{d}+2)<3/2^{d}$. Then we take the difference $F_{\mathrm{th,dp}}-3/2^{d}$ as:
$F_{\mathrm{th,dp}}-\frac{3}{2^{d}}=\frac{(2^{d}-1)\left(2^{d}-2+\sqrt{(2^{d}-2)^{2}+2^{d+3}d}\right)-2^{d+2}d}{2^{2d+1}d},$ (46)
whose positivity is only determined by the numerator. Then notice that $2^{d}-1\geq 2^{d}-2\geq 2^{d-1}$ for $d\geq 2$. Thus we can relax the first product in the numerator:
$(2^{d}-1)\left(2^{d}-2+\sqrt{(2^{d}-2)^{2}+2^{d+3}d}\right)\geq 2^{d-1}(2^{d-1}+2^{d-1})=2^{2d-1}.$ (47)
It is then easy to show that $2^{d-1}>4d$ for $d\geq 6$. Then with straightforward calculation, we can also verify that $F_{\mathrm{th,dp}}>3/(2^{d}+2)$ for $d=2,3,4,5$. ∎
III.1.1 Comment on rank-2 dephased GHZ state
We can also consider the worst-case scenario, i.e. the $d$-qubit rank-2 dephased GHZ state $\rho=F|\mathrm{GHZ}_{d}\rangle\langle\mathrm{GHZ}_{d}|+(1-F)Z|\mathrm{GHZ}_{d}\rangle\langle\mathrm{GHZ}_{d}|Z$. The fidelity threshold for this type of noisy GHZ state to be advantageous over the optimal sensing strategy is:
$F_{\mathrm{th},r=2}=\frac{1+\sqrt{d}}{2\sqrt{d}}>\frac{1}{2}.$ (48)
In fact, it can be easily shown using GME criterion in [67] that any rank-2 GHZ-diagonal state $\rho=F|\mathrm{GHZ}_{d}\rangle\langle\mathrm{GHZ}_{d}|+(1-F)P|\mathrm{GHZ}_{d}\rangle\langle\mathrm{GHZ}_{d}|P$, where $P$ is a $d$-qubit Pauli string that does not stabilize $|\mathrm{GHZ}_{d}\rangle$, is GME regardless of fidelity $F$. Therefore, although the rank-2 dephased GHZ state is always GME, it is not very metrologically useful for our task. However, in reality the prepared probe state is very unlikely to be rank-2.
From another perspective, the results for dephased GHZ states suggesst that GME is not sufficient for demonstrating quantum advantage in DQS. Recall that we have also demonstrated the unnecessity of GME for quantum advantage in DQS with the example of depolarized GHZ states. Consequently, we have rigorously demonstrated that GME is neither sufficient nor necessary for quantum advantage in DQS with concrete examples.
III.2 Imperfect local entanglement generation
We first consider $nd$-qubit depolarized GHZ state with fidelity $F$, and the coefficient $C$ before $n^{2}$ for the second term in QFI derivation is:
$C_{\mathrm{dp}}(F,d,n)=2\left[\frac{2F\frac{1-F}{2^{nd}-1}}{F+\frac{1-F}{2^{nd}-1}}+(2^{nd-1}-1)\frac{2\left(\frac{1-F}{2^{nd}-1}\right)^{2}}{2\frac{1-F}{2^{nd}-1}}\right]=\frac{(1-F)(4^{nd}F-2^{nd}-2)}{[(2^{nd}-2)F+1](2^{nd}-1)}.$ (49)
Similar to perfect local entanglement generation case, we can derive the fidelity threshold for collectively depolarized $nd$-qubit GHZ state to demonstrate advantage in parameter average estimation as:
$F_{\mathrm{th,dp}}(F,d,n)=2^{-nd}+\frac{(2^{nd}-1)\left(2^{nd}-2+\sqrt{(2^{nd}-2)^{2}+2^{nd+3}d}\right)}{2^{2nd+1}d}.$ (50)
The threshold also quickly converges to $1/d$. Then we can express $C_{\mathrm{dp}}$ by substituting $F$ with $F(n)=k^{n-1}F$ in Eq. 49:
$C_{\mathrm{dp}}(F,d,n,k)=\frac{(1-k^{n-1}F)(4^{nd}k^{n-1}F-2^{nd}-2)}{[(2^{nd}-2)k^{n-1}F+1](2^{nd}-1)}.$ (51)
We could easily prove some intuitive properties.
Proposition III.2.
Given the above model of imperfect local entanglement generation, $C_{\mathrm{dp}}$ decreases monotonically as initial fidelity $F$, number of local parameters $d$, and local entanglement generation quality $k$, increase for all $n\geq 1$, if $F,k>2^{-d}$.
Proof.
For instance, we could take its partial derivatives with respect to $F,d,k$:
$\frac{\partial}{\partial F}C_{\mathrm{dp}}=\frac{k^{n}\left[1-(2^{d}k)^{n}\frac{F}{k}\right]\left[(3\times 2^{nd}-2)k+2^{nd}(2^{nd}-2)k^{n}F\right]}{(2^{nd}-1)\left[(2^{nd}-2)k^{n}F+k\right]^{2}},$ (52)
$\frac{\partial}{\partial d}C_{\mathrm{dp}}=\ln 2\frac{2^{nd}(k-k^{n}F)\left[1-(2^{d}k)^{n}\frac{F}{k}\right]\left[(3\times 2^{nd}-4)k^{n}F+k\right]n}{(2^{nd}-1)^{2}\left[(2^{nd}-2)k^{n}F+k\right]^{2}},$ (53)
$\frac{\partial}{\partial k}C_{\mathrm{dp}}=\frac{k^{n-1}F\left[1-(2^{d}k)^{n}\frac{F}{k}\right]\left[(3\times 2^{nd}-2)k+2^{nd}(2^{nd}-2)k^{n}F\right](n-1)}{(2^{nd}-1)\left[(2^{nd}-2)k^{n}F+k\right]^{2}}.$ (54)
Notice that all the above partial derivatives have positive denominators for $n\geq 1$, $F\in(0,1]$, $k\in(0,1)$ and $d\geq 2$. The numerators all contain one term $\left[1-(2^{d}k)^{n}\frac{F}{k}\right]$ while the remaining terms are also positive under the parameter regime of our interest. Therefore, the positivity of the partial derivatives is only determined by the positivity of $\left[1-(2^{d}k)^{n}\frac{F}{k}\right]$. It can be easily determined that as long as $F,k>2^{-d}$ this term is negative for all $n\geq 1$. ∎
We consider the asymptotic limit of the QFI when the local entanglement generation is imperfect, i.e. $k<1$. Explicitly, we consider $\lim_{n\rightarrow\infty}d[1-C_{\mathrm{dp}}(F,d,n,k)]n^{2}$:
$0\leq d[1-C_{\mathrm{dp}}(F,d,n,k)]n^{2}=$ $\frac{d\left(k-2^{nd}k^{n}F\right)}{(2^{nd}-1)k\left[k+(2^{nd}-2)k^{n}F\right]}n^{2}$
$\leq$ $\frac{4^{nd}k^{2n}F^{2}d}{(2^{nd}-1)k\left[k+(2^{nd}-2)k^{n}F\right]}n^{2}$
$=$ $\frac{F^{2}d}{(1-2^{-nd})k\left[k^{1-n}2^{-nd}+(1-2^{1-nd})F\right]}k^{n}n^{2}.$ (55)
Then taking the limit of $n\rightarrow\infty$ gives $\lim_{n\rightarrow\infty}d[1-C_{\mathrm{dp}}(F,d,n,k)]n^{2}=0$. This demonstrates that the Heisenberg scaling breaks down and the QFI vanishes when more sensors are entangled through imperfect local entanglement generation.
For the maximal number of local sensors per node, in the main text we estimate that $n_{\mathrm{max}}(d,F,k)\approx-\ln(dF)/\ln(k)$. We can explicitly take the partial derivatives with respect to $F$ and $k$ to evaluate the sensitivities of $n_{\mathrm{max}}$ to changes in $F$ and $k$.
$S_{F}=\frac{\partial}{\partial F}n_{\mathrm{max}}\approx-\frac{\partial}{\partial F}\frac{\ln(dF)}{\ln(k)}=-\frac{1}{F\ln(k)},$ (56)
$S_{k}=\frac{\partial}{\partial k}n_{\mathrm{max}}\approx-\frac{\partial}{\partial k}\frac{\ln(dF)}{\ln(k)}=\frac{\ln(dF)}{k\ln^{2}(k)}.$ (57)
We can then compare the sensitivities $S_{F}$ and $S_{k}$ by taking their ratio:
$\frac{S_{k}}{S_{F}}\approx-\frac{F\ln(dF)}{k\ln(k)}.$ (58)
We can visualize the behavior of the sensitivity ratio when $d=3$ in Fig. 4. It can be observed in general $S_{k}/S_{F}\gg 1$ for high $k$ which is needed for meaningful local entanglement generation.
Figure 4: Ratio between $n_{\mathrm{max}}$’s sensitivity to $k$ and sensitivity to $F$, i.e. $S_{k}/S_{F}$. For this figure $d=3$ is fixed.
In addition, we consider the scenario where the number of local quantum sensors is fixed. The focus is thus on the impact from local entanglement generation quality $k$ and $d$-qubit GHZ state fidelity $F$, respectively. To this end, we examine the partial derivatives of relative advantage $\eta_{\mathrm{dp}}(F,d,n,k)=d[1-C_{\mathrm{dp}}(F,d,n,k)]$ w.r.t. $k$ and $F$, respectively.
$\frac{\partial}{\partial k}\eta_{\mathrm{dp}}=\frac{Fk^{n-2}d\left(2^{nd}k^{n}F-k\right)\left[\left(3\times 2^{nd}-2\right)k+\left(2^{nd}-2\right)2^{nd}k^{n}F\right](n-1)}{\left(2^{nd}-1\right)\left[\left(2^{nd}-2\right)k^{n}F-k\right]^{2}},$ (59)
$\frac{\partial}{\partial F}\eta_{\mathrm{dp}}=\frac{k^{n-1}d\left(2^{nd}k^{n}F-k\right)\left[\left(3\times 2^{nd}-2\right)k+\left(2^{nd}-2\right)2^{nd}k^{n}F\right]}{\left(2^{nd}-1\right)\left[\left(2^{nd}-2\right)k^{n}F-k\right]^{2}}.$ (60)
The partial derivatives are non-negative when $\left(2^{nd}k^{n}F-k\right)\geq 0$, which is generally satisfied in realistic parameter regimes. Moreover, despite the apparent complicated form of each individual expression, we have a simple relationship for their ratio
$\frac{\partial\eta_{\mathrm{dp}}/\partial k}{\partial\eta_{\mathrm{dp}}/\partial F}=(n-1)\frac{F}{k},$ (61)
which demonstrate that when fixed $n$ is small and when $F$ is in general smaller than $k$, the quantum network entanglement distribution quality could indeed have higher impact on the relative advantage. We demonstrate such effect in Fig. 5. It is shown that curves with identical line style but different colors all overlap, but curves with different line styles are still separable from each other. This reveals that the impact from different GHZ state fidelity $F$ is greater than the local entanglement generation quality $k$, because $n=2$ is small and for the considered parameter values we have $k>F$.
Figure 5: Visualization of $\eta_{\mathrm{dp}}$ as function of the number of sensor nodes $d$, for some choices of initial $d$-qubit GHZ state fidelity $F$ and local entanglement generation quality $k$. We fix the number of local sensors $n=2$ while varying $F$ and $k$. Dashed curves correspond to $F=0.8$; solid curves correspond to $F=0.9$; dot-dashed curves correspond to $F=0.99$. Blue color represents $k=0.9999$; yellow color represents $k=0.999$; green color represents $k=0.99$.
III.3 Comparison with local imperfect GHZ state
We have considered the global sensing strategy with imperfect initial $d$-qubit entanglement across the $d$ sensor nodes, together with imperfect local entanglement generation when extending the $d$-qubit probe state to the $nd$-qubit state. If we take into account imperfect local entanglement generation, it is natural to consider that the comparison baseline, the local strategy, will also need to utilize imperfect local entangled probe state.
We consider a specific local strategy, where each node will utilize $n$-qubit imperfect GHZ state to probe the single local parameter which is encoded through coupling with the $Z$ component of the collective spin. In such single-parameter estimation scenario for each node, the QFI be calculated through [63]:
$F_{\rho_{0},H}=$ $2\sum_{a,b}\frac{(\lambda_{a0}-\lambda_{b0})^{2}}{\lambda_{a0}+\lambda_{b0}}|\langle\psi_{a0}|H|\psi_{b0}\rangle|^{2}$
$=$ $2\sum_{a,b}(\lambda_{a0}+\lambda_{b0})|\langle\psi_{a0}|H|\psi_{b0}\rangle|^{2}-\sum_{a,b}\frac{8\lambda_{a0}\lambda_{b0}}{\lambda_{a0}+\lambda_{b0}}|\langle\psi_{a0}|H|\psi_{b0}\rangle|^{2},$ (62)
where $\lambda_{a0}$ are eigenvalues of local initial probe state $\rho_{0}=\sum_{a}\lambda_{a0}|\psi_{a0}\rangle\langle\psi_{a0}|$, and $H=\frac{1}{2}\sum_{i=0}^{n-1}\sigma^{(i)}_{z}$ is the local generator of the parameter-encoding unitary. Similar to the previous twirling argument, here without loss of generality, we consider local initial probe state to be in the form of GHZ-diagonal state. Then through analysis that is same to the multiparameter QFIM case, we can arrive at the analytical form of QFI for initial GHZ-state diagonal state which is identical to the QFIM in multiparameter case:
$\mathcal{F}=n^{2}-n^{2}\sum_{(a,b)\in\mathcal{S}}\frac{2\lambda_{a0}\lambda_{b0}}{\lambda_{a0}+\lambda_{b0}}=(1-C)n^{2},$ (63)
where $\mathcal{S}$ is again the set of index pairs $(a,b)$ such that $|\psi_{a0}\rangle$ and $|\psi_{b0}\rangle$ are superpositions of the same two computational basis states but with opposite relative phase, and we re-emphasize that for such pairs both $(a,b)$ and $(b,a)$ are included in $\mathcal{S}$.
We focus on the impact of imperfect initial probe state preparation, and thus assume that the optimal measurement on each local node can be performed, that is the QCRB for each sensor node’s local estimation can be achieved. Then according to the propagation of error, we have the estimation error for the average of all local parameters with the local strategy:
$\mathrm{Var}_{\mathrm{local}}(\hat{\theta}_{1})=$ $\sum_{l=1}^{d}\left(\frac{\partial\theta_{1}}{\partial x_{l}}\right)^{2}\mathrm{Var}(\hat{x}_{l})=\frac{1}{n^{2}N}\sum_{l=1}^{d}\frac{1}{d(1-C_{l})},$ (64)
for $N$ repetitions of measurements, $d$ sensor nodes, and $n$ qubits per node.
We consider the following model of imperfect local entanglement generation: The generated local probe state at each node is an identical depolarized GHZ state with fidelity $F(n)=\tilde{k}^{n-1}$ as a function of the number of local sensors $n$, where $\tilde{k}\in(0,1)$ is a constant representing the quality of local entanglement generation and the higher the better. To make the comparison with the global strategy, we consider the following correspondence: $\tilde{k}=\sqrt[d]{k}$, where $k$ is the same constant as we used for the fidelity of the $nd$-qubit global probe state with imperfect local entanglement generation. This correspondence is motivated by the fact that when $n$ increases one, $d$ qubits are added to the global probe state, while only one qubit is added to each node’s local probe state. Given this model, we can further express $\mathrm{Var}_{\mathrm{local}}(\hat{\theta}_{1})$ as:
$\mathrm{Var}_{\mathrm{local}}(\hat{\theta}_{1})=\frac{1}{(1-C_{\mathrm{local}})n^{2}N},$ (65)
where
$C_{\mathrm{local}}(d,n,k)=\frac{\left(1-k^{(n-1)/d}\right)\left[(2^{n}-2)k^{1/d}+4^{n}k^{n/d}\right]}{(2^{n}-1)\left[k^{1/d}+(2^{n}-2)k^{n/d}\right]}.$ (66)
Then we can make the explicit comparison between $\eta_{\mathrm{global}}=d[1-C_{\mathrm{dp}}(F,d,n,k)]$ for the global strategy and $\eta_{\mathrm{local}}=(1-C_{\mathrm{local}})$ for the local strategy. As an example, we visualize the ratio $r=\eta_{\mathrm{global}}/\eta_{\mathrm{local}}$ when $F=0.9$, $d=3$, and $k=0.9999$, in Fig. 6. It can be observed that the threshold of local sensor number for advantage over imperfect local strategy is higher than the threshold for advantage over the optimal local strategy. This is intuitive as the baseline in the former scenario is worse. However, it is important to re-emphasize that the global strategy performance will become worse than the specific local strategy, even if imperfect local entanglement generation is included in the local strategy. This reinforces the fundamental limit in the advantage of global strategy when local entanglement generation is imperfect. In fact, this can be seen analytically through the asymptotic analysis of $\eta_{\mathrm{local}}$ and $\eta_{\mathrm{global}}$ (under the reasonable assumption of $21>k>1/$ 2):
$\eta_{\mathrm{global}}\sim$ $(dF)k^{n-1},$ (67)
$\eta_{\mathrm{local}}\sim$ $\left(k^{1/d}\right)^{n-1},$ (68)
which means that as $n$ increases $\eta_{\mathrm{global}}$ will always drop below $\eta_{\mathrm{local}}$.
Figure 6: Visualization of the comparison between global and local sensing strategies with imperfect local entanglement generation, when $F=0.9$, $d=3$, and $k=0.9999$. The blue solid line denotes $r=\eta_{\mathrm{global}}/\eta_{\mathrm{local}}$, and the yellow dashed line denotes $\eta_{\mathrm{global}}$, which is identical to $\eta_{\mathrm{dp}}$ in the main text. The black dashed line marks the baseline value of 1: If $\eta_{\mathrm{global}}>1$ there is quantum advantage over the optimal local strategy, and if $r>1$ there is advantage of using global strategy over local strategy when local entanglement generation is imperfect for both.
III.4 Justification of the phenomenological model
Here we justify the validity of our phenomenological model of noisy local entanglement generation under the assumption of not too biased errors.
Specifically, we consider noisy CNOT models where there is $F_{\mathrm{CNOT}}$ probability of realizing the noiseless CNOT and $(1-F_{\mathrm{CNOT}})$ probability of applying Pauli noise channel on the involved qubits. When $d(n-1)$ CNOT gates are applied, there will be $F_{\mathrm{CNOT}}^{d(n-1)}\sim k^{n-1}$ probability to obtain the ideal output state, and $(1-F_{\mathrm{CNOT}}^{d(n-1)})\sim(1-k^{n-1})$ probability to have additional Pauli operators applied to the final state. Our initial state is a GHZ-diagonal state with GHZ fidelity $F$. Thus, we have $Fk^{n-1}$ probability of directly obtaining the ideal GHZ state, $(1-F)k^{n-1}$ probability of obtaining mixture of other orthogonal GHZ states, $F(1-k^{n-1})$ probability of applying Pauli error operators on the ideal GHZ state, and $(1-F)(1-k^{n-1})$ probability of applying Pauli error operators on the mixture of other orthogonal GHZ states. Notice that for each GHZ basis state there are $2^{nd}$ out of $4^{nd}$ $nd$-qubits Pauli strings which will transform it into the objective GHZ state. However, if the errors are not too biased, each error operator has similar probability to be applied. The errors which stabilize the objective GHZ state will only contribute roughly $(2^{nd}/4^{nd})(1-k^{n-1})=(1-k^{n-1})/2^{nd}\sim(1/2)^{nd}$ to the fidelity. Similarly, the errors that transform other orthogonal GHZ states to the objective GHZ state will contribute roughly $(2^{nd}/4^{nd})(1-F)(1-k^{n-1})\sim(1/2)^{nd}$. However, in practice $k\sim F_{\mathrm{CNOT}}^{d}$ is significantly large than $1/2$ for good local entanglement generation, which means that $(1/2)^{nd}\ll Fk^{n-1}$. The above thus justifies that the fidelity of $nd$-qubit GHZ state after noisy local entanglement generation can be approximated by $Fk^{n-1}$.
Then we justify that the depolarizing channel is a valid approximation for error models which are not too biased and close to uniform distribution. We consider the explicit calculation of the parameter $C$ for QFI:
$C=1-\sum_{(a,b)\in\mathcal{S}}\frac{(\lambda_{a0}-\lambda_{b0})^{2}}{\lambda_{a0}+\lambda_{b0}},$ (69)
where we have removed the $1/2$ factor before the summation and correspondingly discard double counting in $\mathcal{S}$. For GHZ state that can potentially demonstrate quantum advantage, we know that its fidelity $F>1/d\gg 2^{-nd}$. For error distribution close to unbiased distribution, we have that $\lambda_{a0}+\lambda_{b0}\sim(1-F)2^{1-nd}$, and $|\lambda_{a0}-\lambda_{b0}|\sim f(\lambda_{a0}+\lambda_{b0})\sim f(1-F)2^{1-nd}$, where $f$ characterizes the fluctuation in the error distribution. Then the summation for calculating $C$ can be estimated as
$\sum_{(a,b)\in\mathcal{S}}\frac{(\lambda_{a0}-\lambda_{b0})^{2}}{\lambda_{a0}+\lambda_{b0}}\sim\frac{(F-\lambda)^{2}}{F+\lambda}+(2^{nd-1}-1)\frac{\left[f(1-F)2^{1-nd}\right]^{2}}{(1-F)2^{1-nd}}\sim F+f^{2}(1-F),$ (70)
where $\lambda\sim 2^{-nd}$ denotes the fidelity of $(|0\dots 0\rangle-|1\dots 1\rangle)/\sqrt{2}$. For unbiased error distributions, we have $f\ll 1$, so that the second term on the second line in the above is negligible, in comparison to the first term which estimates the contribution from $(|0\dots 0\rangle+|1\dots 1\rangle)/\sqrt{2}$. We can visualize the the ratio between the estimated contribution to the summation for $C$ from $(|0\dots 0\rangle+|1\dots 1\rangle)/\sqrt{2}$ and other errors terms, i.e. $f^{2}(1-F)/F$, in Fig. 7. It is clear that the contribution from error components in the GHZ-diagonal state is negligible when the fluctuation in the error distribution is small, i.e. when the error model is not too biased. This thus justifies that the depolarized error quantitatively captures the key impact of unbiased error distributions on the evaluation of relative quantum advantage.
Figure 7: Visualization of the ratio between the estimated contribution from the standard GHZ component and other error terms, $f^{2}(1-F)/F$ as function of GHZ fidelity $F$ for error distribution fluctuation factors $f=0.1,0.01,0.001$.
IV Local measurement for entangled sensors
According to propagation of error, we can obtain the variance of estimating a certain parameter encoded in a quantum state from the measurement results of an observable $M$ [63]:
$\mathrm{Var}_{M}(\hat{\theta}_{1})=\frac{\langle M^{2}\rangle-\langle M\rangle^{2}}{\left|\frac{\partial}{\partial\theta_{1}}\langle M\rangle\right|^{2}},$ (71)
where the expectation value $\langle M\rangle$ is taken under the encoded state $\rho_{x}$ and thus is a function of local parameters $x_{i}$. We note that in quantum sensing, the values of parameters to estimate are usually small. Therefore, when using the propagation of error formula we may take the limit of $\theta\rightarrow 0$. It is straightforward to use the aforementioned orthogonal transformation to convert $x_{i}$ into linear combinations of derived parameters that include the parameter of interest $\theta_{1}$. Gram-Schmidt process can be used to construct orthonormal basis which includes $v_{1}=(1,\dots,1)^{T}/\sqrt{d}$. For instance, for $d=3$ we can construct an orthonormal basis:
$v_{1}=(1,1,1)^{T}/\sqrt{3},$ (72)
$v_{2}=(-1,2,-1)^{T}/\sqrt{6},$ (73)
$v_{3}=(-1,0,1)^{T}/\sqrt{2},$ (74)
where $\theta_{i}=v_{i}^{T}x$, and in our scenario we are only interested in $\theta_{1}$. Given a specific $d$-dimensional orthonormal basis, we will be able to transform the $x$-dependence of $\langle M\rangle$ into $\theta$-dependence. Then the partial derivative with respect to $\theta_{1}$ will be straightforward.
IV.1 Optimal measurement for noiseless case is useless for noisy case
Recall that here the observable of our interest is $M=\sigma_{x}^{\otimes nd}$. For GHZ-diagonal states in general, the expectation value of observable is the average of expectation value under pure GHZ states weighted by the diagonal elements of density matrix in GHZ basis. Since $M^{2}=I^{\otimes nd}$ we always have $\langle M^{2}\rangle=1$. Consider GHZ states corresponding to the same binary string but with opposite relative phase $|\mathrm{GHZ}_{nd}^{\pm}(b)\rangle=(|b\rangle\pm|2^{nd}-b-1\rangle)/\sqrt{2}$, where $|x\rangle$ denote the computational basis state corresponding to the binary representation of integer $x=0,1,\dots,2^{nd}-1$ and $b=0,1,\dots,2^{nd-1}-1$. We have that:
$\langle\mathrm{GHZ}_{nd}^{+}(b)|U^{\dagger}(x)MU(x)|\mathrm{GHZ}_{nd}^{+}(b)\rangle=$ $\frac{1}{2}\left(\langle b|+e^{-i\phi(x)}\langle 2^{nd}-b-1|\right)\sigma_{x}^{\otimes nd}\left(|b\rangle+e^{i\phi(x)}|2^{nd}-b-1\rangle\right)$
$=$ $\frac{1}{2}\left(e^{i\phi(x)}\langle b|\sigma_{x}^{\otimes nd}|2^{nd}-b-1\rangle+e^{-i\phi(x)}\langle 2^{nd}-b-1|\sigma_{x}^{\otimes nd}|b\rangle\right)$
$=$ $\frac{1}{2}\left(e^{i\phi(x)}+e^{-i\phi(x)}\right)=\cos\left(\phi(x)\right)$
$=$ $-\langle\mathrm{GHZ}_{nd}^{-}(b)|U^{\dagger}(x)MU(x)|\mathrm{GHZ}_{nd}^{-}(b)\rangle$ (75)
Given the above properties of GHZ-diagonal state under the encoding channel and observable of our interest, the analysis of depolarized GHZ state becomes significantly simplified. Only $(|0\dots 0\rangle\pm|1\dots 1\rangle)/\sqrt{2}$ contribute to the expectation value $\langle M\rangle$, because other GHZ states corresponding to the same binary string have identical weights and thus their contributions cancel each other. Thus for depolarization error model we have:
$\langle M\rangle_{n}(\theta_{1})=\left(F-\frac{1-F}{2^{nd}-1}\right)\cos\left(n\sqrt{d}\theta_{1}\right),$ (76)
where $n$ denotes the number of sensors per node, and noiseless local entanglement generation corresponds to $n=1$. Then we can substitute the above into Eq. 71:
$\mathrm{Var}_{M}(\hat{\theta}_{1})=\frac{1-\left(F-\frac{1-F}{2^{nd}-1}\right)^{2}\cos^{2}\left(n\sqrt{d}\theta_{1}\right)}{dn^{2}\left(F-\frac{1-F}{2^{nd}-1}\right)^{2}\sin^{2}\left(n\sqrt{d}\theta_{1}\right)}$ (77)
It is obvious that when taking the above function to the limit of $\theta_{1}\rightarrow 0$ it goes to infinity if $F<1$, which suggests that this specific measurement scheme is useless to estimate small values with high accuracy. We comment that the divergence of estimation variance in small local parameter regime is general for any GHZ-diagonal state with fidelity below 1, because the denominator in error propagation formula will approach zero, while the numerator will stay non-zero when the fidelity is not equal to 1.
IV.2 Optimization of restricted local measurement
Figure 8: The estimation variance as a function of azimuthal angle $\alpha$ is shown as the blue solid curve, for $F=0.8,d=3,n=1$. The black dashed line denotes the baseline which can be achieved by the best local strategy. The yellow solid line demonstrates the inverse of the QFI, representing the ultimate estimation variance which requires some other measurement.
The space of local measurement schemes is large. For concreteness and simplicity, here we focus on a specific family of measurement characterized by one parameter $\alpha$: $M(\alpha)=\left[O(\alpha)\right]^{\otimes nd}$ where $O(\alpha)=|\psi^{+}(\alpha)\rangle\langle\psi^{+}(\alpha)|-|\psi^{-}(\alpha)\rangle\langle\psi^{-}(\alpha)|$ with $|\psi^{\pm}(\alpha)\rangle=(|0\rangle\pm e^{i\alpha}|1\rangle)$. That is, $O(\alpha)=e^{i\alpha}|1\rangle\langle 0|+e^{-i\alpha}|0\rangle\langle 1|$ while the optimal measurement in noiseless case is $M(0)$. Notice that $\left(O(\alpha)\right)^{2}=I$ and $O(\alpha)$ is always off-diagonal in computational basis. Thus we can still simplify its expectation values under GHZ-diagonal states, especially depolarized GHZ states with $nd$ qubits:
$\mathrm{Var}_{M(\alpha)}(\hat{\theta}_{1})=\frac{1-\left(F-\frac{1-F}{2^{nd}-1}\right)^{2}\cos^{2}\left[n\sqrt{d}(\theta_{1}+\sqrt{d}\alpha)\right]}{dn^{2}\left(F-\frac{1-F}{2^{nd}-1}\right)^{2}\sin^{2}\left[n\sqrt{d}(\theta_{1}+\sqrt{d}\alpha)\right]}.$ (78)
It is clear that the non-zero azimuthal angle $\alpha$ has the effect of modifying the undesired zero denominator when $\theta_{1}=0$. Then we can safely take $\theta_{1}=0$ and optimize $\alpha$:
$\left.\mathrm{Var}_{M(\alpha)}(\hat{\theta}_{1})\right|_{\theta_{1}=0}=\frac{1-\left(F-\frac{1-F}{2^{nd}-1}\right)^{2}\cos^{2}\left(nd\alpha\right)}{dn^{2}\left(F-\frac{1-F}{2^{nd}-1}\right)^{2}\sin^{2}\left(nd\alpha\right)}.$ (79)
The partial derivative with respect to $\alpha$ gives rise to:
$\frac{\partial}{\partial\alpha}\left.\mathrm{Var}_{M(\alpha)}(\hat{\theta}_{1})\right|_{\theta_{1}=0}=-\frac{2^{nd+1}(1-F)\left[2^{nd}(1+F)-2\right]}{n\left(2^{nd}F-1\right)^{2}}\frac{\cos(nd\alpha)}{\sin^{3}(nd\alpha)}.$ (80)
It is thus obvious that the variance takes the minimum values when $\cos(nd\alpha_{\mathrm{opt}})=0$, i.e. $\alpha_{\mathrm{opt}}=\frac{2l+1}{2nd}\pi$ with $l\in\mathbb{Z}$. The minimum value is:
$\left.\mathrm{Var}_{M(\alpha_{\mathrm{opt}})}(\hat{\theta}_{1})\right|_{\theta_{1}=0}=\frac{1}{d\left(F-\frac{1-F}{2^{nd}-1}\right)^{2}n^{2}}=\frac{1}{\eta_{M(\alpha_{\mathrm{opt}})}n^{2}}.$ (81)
Such periodicity of the optimal azimuthal angle echoes with previous numerical results [70]. We can visualize the parameter estimation variance with different $\alpha$ for $n=1$ in Fig. 8. The ratio between $\eta_{M(\alpha_{\mathrm{opt}})}$ and $\eta_{\mathrm{dp}}(k=1)$ is:
$\frac{\eta_{M(\alpha_{\mathrm{opt}})}}{\eta_{\mathrm{dp}}(k=1)}=F+\frac{1-F}{2^{nd}-1},$ (82)
which quickly converges to $F$ for larger $n$ and $d$.
The fidelity threshold for $nd$-qubit depolarized GHZ state to be advantageous over the optimal local strategy when using the optimized azimuthal measurement is given by $\eta_{M(\alpha_{\mathrm{opt}})}=1$:
$F_{\mathrm{th},M(\alpha_{\mathrm{opt}})}(n)=\frac{2^{nd}+\sqrt{d}-1}{2^{nd}\sqrt{d}}.$ (83)
For $n=1$, we have the fidelity threshold for the initial depolarized GHZ state to demonstrate advantage when using the optimized azimuthal measurement, i.e. $\eta_{M(\alpha_{\mathrm{opt}})}>1$:
$F_{\mathrm{th},M(\alpha_{\mathrm{opt}})}(1)=\frac{2^{d}+\sqrt{d}-1}{2^{d}\sqrt{d}},$ (84)
which can be easily shown to be monotonically decreasing for the distributed regime $d\geq 2$ of our interest.
We can further characterize the increase of estimation variance when the azimuthal angle of the measurement deviates from the optimal value. Firstly, we can explicitly characterize the behavior of estimation variance of the local azimuthal measurement when the azimuthal angle has small deviation from the optimal angle: $\alpha=\alpha_{\mathrm{opt}}+\delta$
$\left.\mathrm{Var}_{M(\alpha_{\mathrm{opt}}+\delta)}(\hat{\theta}_{1})\right|_{\theta_{1}=0}=\frac{1}{d\left(F-\frac{1-F}{2^{nd}-1}\right)^{2}n^{2}}+\frac{d(1-F)2^{nd}\left[2^{nd}(F+1)-2\right]}{\left(2^{nd}F-1\right)^{2}}\delta^{2}+O(\delta^{4}).$ (85)
Then we examine the range of azimuthal angle within which we can achieve better estimation variance than the optimal local sensing strategy. Specifically, we let the estimation variance of azimuthal local measurement in Eq. 79 equal to the estimation variance of the optimal local sensing strategy, i.e.
$\frac{1-\left(F-\frac{1-F}{2^{nd}-1}\right)^{2}\cos^{2}\left(nd\alpha\right)}{dn^{2}\left(F-\frac{1-F}{2^{nd}-1}\right)^{2}\sin^{2}\left(nd\alpha\right)}=\frac{1}{n^{2}}\implies\frac{1-\left(F-\frac{1-F}{2^{nd}-1}\right)^{2}\cos^{2}\left(nd\alpha\right)}{d\left(F-\frac{1-F}{2^{nd}-1}\right)^{2}\sin^{2}\left(nd\alpha\right)}=1,$ (86)
which can be simplified as
$\sin^{2}\left(nd\alpha\right)=\frac{1-\left(F-\frac{1-F}{2^{nd}-1}\right)^{2}}{(d-1)\left(F-\frac{1-F}{2^{nd}-1}\right)^{2}}.$ (87)
The solution to the above equation determines the azimuthal angles at which the estimation variance of local azimuthal measurement equals the estimation variance of the optimal local sensing strategy. As the estimation variance of local azimuthal measurement has a period of $T=\pi/nd$, the range of the azimuthal angle for DQS quantum advantage is
$W_{\alpha}=\frac{\pi}{nd}-\frac{2}{nd}\arcsin\sqrt{\frac{1-\left(F-\frac{1-F}{2^{nd}-1}\right)^{2}}{(d-1)\left(F-\frac{1-F}{2^{nd}-1}\right)^{2}}}.$ (88)
This range of allowed azimuthal angle can be considered as an indicator of the robustness of DQS quantum advantage to local quantum control for azimuthal measurement. To characterize the properties of $W_{\alpha}$, we further consider the ratio between the range which allows quantum advantage $W_{\alpha}$ and the period $T$, $R_{\alpha}=W_{\alpha}/T$,
$R_{\alpha}(F,n,d)=1-\frac{2}{\pi}\arcsin\sqrt{\frac{1-\left(F-\frac{1-F}{2^{nd}-1}\right)^{2}}{(d-1)\left(F-\frac{1-F}{2^{nd}-1}\right)^{2}}}.$ (89)
We have the following basic properties of $R_{\alpha}(F,n,d)$.
Proposition IV.1.
$R_{\alpha}(F,n,d)$ increases monotonically as the fidelity $F$, the number of local quantum sensors $n$, and the number of sensor nodes $d$ increase.
Proof.
The properties can be proved through explicit evaluation of the partial derivatives. We first take the partial derivative w.r.t. the fidelity $F$
$\frac{\partial}{\partial F}R_{\alpha}=\frac{2^{\frac{nd}{2}+1}\left(2^{nd}-1\right)^{2}}{\pi\left(2^{nd}F-1\right)\sqrt{(1-F)\left[2^{nd}(1+F)-2\right]\left[d\left(2^{nd}F-1\right)^{2}-\left(2^{nd}-1\right)^{2}\right]}}\geq 0.$ (90)
Then we take the partial derivative w.r.t. the number of local quantum sensors $n$
$\frac{\partial}{\partial n}R_{\alpha}=\frac{\ln(2)(1-F)2^{\frac{nd}{2}+1}\left(2^{nd}-1\right)d}{\pi\left(2^{nd}F-1\right)\sqrt{(1-F)\left[2^{nd}(1+F)-2\right]\left[d\left(2^{nd}F-1\right)^{2}-\left(2^{nd}-1\right)^{2}\right]}}\geq 0.$ (91)
Lastly we take the partial derivative w.r.t. the number of sensor nodes $d$
$\frac{\partial}{\partial n}R_{\alpha}=\frac{2^{\frac{nd}{2}}\left[4^{nd}F(1-F^{2})-2^{nd}(1-F)(1+3F+(d-1)n\ln(4))-(1-F)((d-1)n\ln(4)-2)\right]}{\pi(d-1)\left(2^{nd}F-1\right)\sqrt{(1-F)\left[2^{nd}(1+F)-2\right]}\left[d\left(2^{nd}F-1\right)^{2}-\left(2^{nd}-1\right)^{2}\right]}\geq 0.$ (92)
For the above inequality we consider the last two terms in the numerator
$-2^{nd}(1-F)(1+3F+(d-1)n\ln(4))-(1-F)((d-1)n\ln(4)-2)$
$=$ $-(1-F)\left[\left(2^{nd}(1+3F)-2\right)-\left(2^{nd}-1\right)(d-1)n\ln(4)\right]$
$\geq$ $-(1-F)\left[4\times 2^{nd}-\left(2^{nd}-1\right)\right]=-(1-F)\left(3\times 2^{nd}+1\right),$ (93)
where for the inequality we have used $F\leq 1$, $d\geq 2$ in DQS setup, and $n\geq 1$. Then we evaluate the lower bound of the entire numerator
$4^{nd}F(1-F^{2})-2^{nd}(1-F)(1+3F+(d-1)n\ln(4))-(1-F)((d-1)n\ln(4)-2)$
$\geq$ $4^{nd}F(1-F^{2})-(1-F)\left(3\times 2^{nd}+1\right)$
$=$ $(1-F)\left[4^{nd}F(1+F)-\left(3\times 2^{nd}+1\right)\right]$
$\geq$ $(1-F)\left[4^{nd}\frac{1}{\sqrt{d}}\left(1+\frac{1}{\sqrt{d}}\right)-4\times 2^{nd}\right]\geq 0,$ (94)
where for the second inequality we have used the fidelity threshold for DQS advantage using the optimized azimuthal measurement: We know that if fidelity is below the threshold we have $R_{\alpha}=0$. ∎
Although $R_{\alpha}(F,n,d)$ is monotonically increasing, it quickly converges to a value independent of $n$:
$R_{\alpha}(F,n,d)\to 1-\frac{2}{\pi}\arcsin\sqrt{\frac{1-F^{2}}{(d-1)F^{2}}}.$ (95)
Then we are clear about the behavior of $W_{\alpha}$ as $n$ and $d$ increases for larger relative quantum advantage in DQS
$W_{\alpha}\to\frac{\pi}{nd}\left[1-\frac{2}{\pi}\arcsin\sqrt{\frac{1-F^{2}}{(d-1)F^{2}}}\right].$ (96)
Such scaling clearly demonstrates that while when $n$ and $d$ increase we can in principle achieve higher quantum advantage, the requirement on control also becomes stricter.
We re-emphasize that our optimization of the azimuthal measurement is not a global optimization, as it is locally separable when there are multiple sensor qubits per node. In principle we could utilize entangling measurement locally. However, the fact that when $n=1$ the optimized azimuthal measurement could not saturate the QCRB implies that the QCRB for this distributed quantum sensing problem is not achievable with local measurement. It is known that the QCRB can be achieved by projective measurement, and when $n=1$ the local projective measurement should be a tensor product of single-qubit projective measurements. Moreover, the assumed depolarization noise on the GHZ state guarantees symmetry among all qubits, so it suffices to consider identical projective measurement per qubit. Note that under our problem setup, z-direction projection is unable to extract the information of the parameter to estimate. Therefore, our optimization of azimuthal projective measurement should have covered the optimal local measurement, while as seen in the results they could not achieve the QCRB. On the other hand, suppose we insist on applying the global entangling measurement, it will need distribution of additional entangled state by the quantum network as resource to implement the measurement. Consequently, the performance of parameter estimation will be further limited by the fidelity of resource state for performing entangling measurement, and thus the fidelity of distributed entangled state by the network must be high. However, as we have seen that when the network is able to distributed high fidelity entangled state, performing local optimized azimuthal measurement can already achieve fairly low estimation variance, which is only $\sim 1/F$ times the ultimate achievable variance by the globally optimal measurement.
IV.2.1 Comment on rank-2 dephased GHZ state
We can also consider the azimuthal measurement for the $d$-qubit, i.e. $n=1$, rank-2 dephased GHZ state $\rho=F|\mathrm{GHZ}_{d}\rangle\langle\mathrm{GHZ}_{d}|+(1-F)Z|\mathrm{GHZ}_{d}\rangle\langle\mathrm{GHZ}_{d}|Z$. Then in the small parameter regime $\theta_{1}\rightarrow 0$, the parameter estimation variance as a function of the azimuthal angle $\alpha$ is:
$\left.\mathrm{Var}_{M(\alpha),r=2}(\hat{\theta}_{1})\right|_{\theta_{1}=0}=\frac{1-\left(2F-1\right)^{2}\cos^{2}\left(d\alpha\right)}{d\left(2F-1\right)^{2}\sin^{2}\left(d\alpha\right)}.$ (97)
Then we can take the partial derivative with respect to $\alpha$ to perform optimization:
$\frac{\partial}{\partial\alpha}\left.\mathrm{Var}_{M(\alpha),r=2}(\hat{\theta}_{1})\right|_{\theta_{1}=0}=-\frac{8F(1-F)}{(2F-1)^{2}}\frac{\cos(d\alpha)}{\sin^{3}(d\alpha)}.$ (98)
Therefore, the lowest variance is achieved when $\cos(d\alpha_{\mathrm{opt}})=0$, i.e. $\alpha_{\mathrm{opt}}=\frac{2l+1}{2d}\pi$ with $l\in\mathbb{Z}$, and this condition is exactly the same for the depolarized GHZ state with $n=1$. Then the minimal variance is:
$\left.\mathrm{Var}_{M(\alpha_{\mathrm{opt}}),r=2}(\hat{\theta}_{1})\right|_{\theta_{1}=0}=\frac{1}{d(2F-1)^{2}}=\frac{1}{\mathcal{F}}.$ (99)
This result means that the optimal azimuthal measurement is able to saturate the QCRB for rank-2 dephased GHZ state.
IV.3 Bell state fidelity requirement estimation
Moreover, we can estimate the fidelity requirement of bipartite entanglement (Bell pair) distribution network to achieve quantum advantage in the local parameter average estimation task, by taking the $(d-1)$-th root of the fidelity threshold. This estimation comes from the assumption that we need $(d-1)$ bipartite entangled states between the sensor nodes to assemble the desired $d$-qubit GHZ state, and the approximation that the final GHZ state has fidelity equal to the product of all Bell states’ fidelities. Specifically, for $n=1$ we have:
$F_{\mathrm{th},M(\alpha_{\mathrm{opt}})}^{\mathrm{Bell}}=\left(\frac{2^{d}-1+\sqrt{d}}{2^{d}\sqrt{d}}\right)^{\frac{1}{d-1}},$ (100)
$F_{\mathrm{th,opt}}^{\mathrm{Bell}}=\left(2^{-d}+\frac{(2^{d}-1)\left(2^{d}-2+\sqrt{(2^{d}-2)^{2}+2^{d+3}d}\right)}{2^{2d+1}d}\right)^{\frac{1}{d-1}},$ (101)
where we use the superscript “Bell” to emphasize that the above fidelity thresholds are for Bell states distributed by the quantum network. The two thresholds are visualized in Fig. 9(a). It can be seen that the fidelity threshold for the global optimal measurement slightly decreases when the number of sensor nodes $d$ is small. More specifically, we have:
$F_{\mathrm{th,opt}}^{\mathrm{Bell}}(2)\approx 0.730,~{}F_{\mathrm{th,opt}}^{\mathrm{Bell}}(3)\approx 0.714,~{}F_{\mathrm{th,opt}}^{\mathrm{Bell}}(4)\approx 0.711,$
$F_{\mathrm{th,opt}}^{\mathrm{Bell}}(5)\approx 0.716,~{}F_{\mathrm{th,opt}}^{\mathrm{Bell}}(6)\approx 0.726,~{}F_{\mathrm{th,opt}}^{\mathrm{Bell}}(7)\approx 0.738.$ (102)
Nevertheless, in general both the thresholds increase monotonically as the number of sensor nodes increases. Moreover, we can straightforwardly evaluate the asymptotic scaling of both thresholds:
$F_{\mathrm{th},M(\alpha_{\mathrm{opt}})}^{\mathrm{Bell}}\sim$ $d^{\frac{1}{2(1-d)}},$ (103)
$F_{\mathrm{th,opt}}^{\mathrm{Bell}}\sim$ $d^{\frac{1}{1-d}}.$ (104)
Very interestingly, the Bell state fidelity threshold for the global optimal measurement is the square of the threshold for the optimized azimuthal measurement in the asymptotic limit of $d\rightarrow\infty$, i.e. $F_{\mathrm{th,opt}}^{\mathrm{Bell}}\sim\left(F_{\mathrm{th},M(\alpha_{\mathrm{opt}})}^{\mathrm{Bell}}\right)^{2}$. It is then easily seen that both fidelity thresholds converge to one in the large $d$ limit. To visualize the asymptotic scaling, we further plot the thresholds of infidelity $\epsilon=1-F$ in log-log coordinate in Fig. 9(b). It is clear that $\epsilon_{\mathrm{th,opt}}^{\mathrm{Bell}}/2$ is almost equal to $\epsilon_{\mathrm{th},M(\alpha_{\mathrm{opt}})}^{\mathrm{Bell}}$ when $d$ becomes large, which verifies the quadratic relation of the asymptotic scalings.
Figure 9: Estimation of Bell state fidelity thresholds for demonstrating quantum advantage in distributed quantum sensing, where the initial $d$-qubit GHZ state is assembled from $(d-1)$ Bell states distributed by the quantum network. (a) The Bell state fidelity threshold when the global optimal measurement is allowed is depicted by the blue curve, while the threshold when only the optimized azimuthal measurement can be performed is illustrated by the red curve. (b) The infidelity $\epsilon=1-F$ thresholds were plotted in the log-log coordinate for additional visualization insight. The blue curve again corresponds to the global optimal measurement and the red curve to the optimized azimuthal measurement. For reference, we plot the half of the infidelity threshold for the global optimal measurement in black dashed line, to better demonstrate the relationship between the asymptotic scalings of both thresholds.
V Parameter estimation with failure in entanglement distribution
In practical quantum networks, there is always non-zero probability to fail in the generation of the demanded entangled state. In our multiparameter estimation scenario where we want to estimate the average of all local parameters, and the probe state we want is a $d$-qubit GHZ state that entangles all $d$ sensor nodes. The state is generated through assembling bipartite entangled states distributed by the quantum network. What might happen is that some bipartite entangled states are not successfully generated within certain attempts.
V.1 Hybrid strategy
We consider a specific way of bipartite entangled state assembly to generate the GHZ state: We assume there is a center node, which will share bipartite entanglement with other nodes. Therefore, for the nodes which fail to establish entangled link with the center node, they will remain separable from other nodes, while all the nodes that successfully share entangled link with the center node will be entangled together. Let $\mathcal{N}$ s.t. $|\mathcal{N}|=d$ be the index set of all sensor noes, and $\mathcal{C}$ be the index set for the nodes which remain isolated, thus the set difference $\mathcal{N}\setminus\mathcal{C}$ denotes the index set of nodes that will be entangled. For the objective parameter to estimate $\theta_{1}=\sum_{i\in\mathcal{N}}x_{i}/\sqrt{d}$, it can be rewritten as:
$\theta_{1}=$ $\frac{1}{\sqrt{d}}\sum_{i\in\mathcal{C}}x_{i}+\sqrt{\frac{|\mathcal{N}\setminus\mathcal{C}|}{d}}\left(\frac{1}{\sqrt{|\mathcal{N}\setminus\mathcal{C}|}}\sum_{j\in\mathcal{N}\setminus\mathcal{C}}x_{j}\right)$
$=$ $\frac{1}{\sqrt{d}}\sum_{i\in\mathcal{C}}x_{i}+\sqrt{\frac{|\mathcal{N}\setminus\mathcal{C}|}{d}}\theta^{\prime}_{1},$ (105)
where the second term is proportional to the average of local parameters on all the nodes that are entangled, $\theta^{\prime}_{1}$. When not all sensor nodes are entangled, we consider that the sensor network will use the following hybrid strategy: The isolated nodes use local probe state to estimate the local parameters $x_{i},i\in\mathcal{C}$ individually, while the entangled nodes use a globally entangled probe state to estimate $\theta^{\prime}_{1}$. Then according to propagation of errors, we have the variance of estimating $\theta$ by such a hybrid strategy as:
$\mathrm{Var}_{\mathrm{hybrid}}(\hat{\theta}_{1,\mathcal{C}})=\frac{1}{d}\sum_{i\in\mathcal{C}}\mathrm{Var}(\hat{x}_{i})+\frac{|\mathcal{N}\setminus\mathcal{C}|}{d}\mathrm{Var}(\hat{\theta}^{\prime}_{1}),$ (106)
where the subscript $\mathcal{C}$ for estimator $\hat{\theta}_{1,\mathcal{C}}$ emphasizes that the estimator is uniquely determined by $\mathcal{C}$. We may also call $\mathcal{C}$ a configuration.
V.2 Combining different configurations
Due to the probabilistic nature of remote entanglement distribution over the quantum repeater networks under our consideration, for each attempt of probe state generation the configuration $\mathcal{C}$ can be different. Therefore, when we repeat the quantum sensing cycles for many shots to accumulate statistical data, the data may correspond to different configurations, and moreover, we know the correspondence exactly due to the heralded nature of quantum repeater network protocols. Let $\mathfrak{C}$ denotes the collection of all possible configurations $\mathcal{C}$. For the scenario with $d$ sensor nodes, we have that $|\mathfrak{C}|=\sum_{n=0}^{d-2}\binom{d}{n}+1=2^{d}-d$ without over-counting, because when $|\mathcal{C}|=d-1$ it is equivalent to that all nodes are isolated.
We would like to utilize all data to increase the estimation accuracy. Suppose we repeat the quantum sensing cycle for $N$ times so that we have $N$ data points, and each configuration $\mathcal{C}$ occurs with probability $p_{\mathcal{C}}$. That is, we may expect that there are $N_{\mathcal{C}}=p_{\mathcal{C}}N$ data points corresponding to configuration $\mathcal{C}$, then we have $\mathrm{Var}(\hat{\theta}_{1,\mathcal{C}})\propto 1/N_{\mathcal{C}}$. Given the assumption of (locally) unbiased estimators, we know that the normalized linear combination of $\hat{\theta}_{1,\mathcal{C}}$ still has the mean of $\theta_{1}$. Then our objective is to minimize the variance of the normalized linear combination:
$\hat{\theta}_{1}=\sum_{\mathcal{C}\in\mathfrak{C}}w_{\mathcal{C}}\hat{\theta}_{1,\mathcal{C}},~{}\mathrm{s.t.}~{}\sum_{\mathcal{C}}w_{\mathcal{C}}=1.$ (107)
We may further assume that $\hat{\theta}_{1,\mathcal{C}}$ are uncorrelated, which according to the Bienaymé’s identity gives us:
$\mathrm{Var}(\hat{\theta}_{1})=\sum_{\mathcal{C}\in\mathfrak{C}}w_{\mathcal{C}}^{2}\mathrm{Var}(\hat{\theta}_{1,\mathcal{C}}).$ (108)
Then it can be derived using Lagrange multiplier that the optimal weighting for the minimum variance is:
$w_{\mathcal{C}}=\frac{1}{\mathrm{Var}(\hat{\theta}_{1,\mathcal{C}})}\left[\sum_{\mathcal{C}\in\mathfrak{C}}\frac{1}{\mathrm{Var}(\hat{\theta}_{1,\mathcal{C}})}\right]^{-1},$ (109)
which is the so-called inverse-variance weighting that gives the minimum variance:
$\mathrm{Var}_{\mathrm{min}}(\hat{\theta}_{1})=$ $\sum_{\mathcal{C}\in\mathfrak{C}}\frac{1}{\mathrm{Var}^{2}(\hat{\theta}_{1,\mathcal{C}})}\left[\sum_{\mathcal{C}\in\mathfrak{C}}\frac{1}{\mathrm{Var}(\hat{\theta}_{1,\mathcal{C}})}\right]^{-2}\mathrm{Var}(\hat{\theta}_{1,\mathcal{C}})=\left[\sum_{\mathcal{C}\in\mathfrak{C}}\frac{1}{\mathrm{Var}(\hat{\theta}_{1,\mathcal{C}})}\right]^{-1}.$ (110)
We comment that when the problem scale $d$ increases, the size of configuration space $\mathfrak{C}$ increases exponentially. Therefore, the minimization of estimation variance by combining the data from all possible configurations will become practically impossible eventually if the problem scale is large. However, when the quantum repeater network is low loss and low error, it is almost guaranteed that every attempt of probe state generation will succeed with a $d$-qubit GHZ state. In such cases, it is good enough to use only the data points which correspond to a complete $d$-qubit GHZ state to estimate $\theta_{1}$.
On the other hand, we may coarse grain the configurations to account only the number of nodes which are entangled for approximation. Thus the corresponding GHZ states are the ensemble average of the GHZ states in different configurations. Let $\mathfrak{C}_{n}$ denote the collection of $\mathcal{C}$ s.t. $|\mathcal{C}|=n$, where $n=0,1,\dots,d-2,d$. Specifically, we consider that $\mathrm{Var}(\hat{\theta}_{1,\mathcal{C}})=F_{\mathcal{C}}/N_{\mathcal{C}}$. In this way, we have the approximate minimum variance:
$\mathrm{Var}_{\mathrm{min}}(\hat{\theta}_{1})=$ $\left[\sum_{\mathcal{C}\in\mathfrak{C}}\frac{1}{\mathrm{Var}(\hat{\theta}_{1,\mathcal{C}})}\right]^{-1}=\left[\sum_{\mathcal{C}\in\mathfrak{C}}\frac{N_{\mathcal{C}}}{F_{\mathcal{C}}}\right]^{-1}$
$=$ $\left[\frac{N_{d}}{F_{d}}+\sum_{n=0}^{d-2}\sum_{\mathcal{C}\in\mathfrak{C}_{n}}\frac{N_{\mathcal{C}}}{F_{\mathcal{C}}}\right]^{-1}$
$\approx$ $\left[\frac{N_{d}}{F_{d}}+\sum_{n=0}^{d-2}\frac{\sum_{\mathcal{C}\in\mathfrak{C}_{n}}N_{\mathcal{C}}}{F_{n}}\right]^{-1}=\left[\frac{N_{d}}{F_{d}}+\sum_{n=0}^{d-2}\frac{N_{n}}{F_{n}}\right]^{-1},$ (111)
where $F_{n}$ denotes the coefficient of variance for configurations with $n$ isolated nodes, and $N_{n}$ denotes the total number of data points for configurations with $n$ isolated nodes.
VI Probe state assembly from bipartite entanglement for DQS
Figure 10: Circuits of (a) CNOT teleportation and (b) GHZ merging. The standard GHZ generation circuit is included in the GHZ merging circuit as the first part.
The preparation of the initial probe state for DQS is envisioned to be based on bipartite entanglement distributed between sensor nodes by quantum networks. Sensor nodes will perform local operations and classical communication (LOCC) to create multipartite entangled state across themselves as the probe state. In this section, we describe two methods to do so, namely gate teleportation and GHZ merging. Before going into details, we comment that in practice the quantum networks might be hybrid, in that different functions are realized by different physical systems. For instance, communication qubits which are in charge of generating and distributing entanglement might be different from the quantum sensors used for DQS. Therefore, to make DQS a reality, experimental development of quantum interconnects [128, 129] is indispensable.
VI.1 Gate teleportation
A Bell pair can be used to perform CNOT teleportation, and the circuit of CNOT teleportation is shown in Fig. 10(a). Then we can directly run the standard GHZ generation circuit, assuming the availability of two-qubit gates between communication qubits and sensor qubits. The standard GHZ generation circuit is shown in the first part of Fig. 10(b).
We can estimate the resources needed for this approach. All generated Bell pairs are consumed, and thus to generate probe state of size $N$, $N$ other sensor qubits (one of them is the center) need to be initialized when performing CNOT teleportation. According to GHZ state generation circuit, in total $(N-1)$ CNOTs are needed, equal to $(N-1)$ Bell pairs. During each gate teleportation after two single qubit measurements, (in ideal case) there are 1/4 probability that no local unitary correction is needed, 1/2 probability that one local unitary correction is needed, and 1/4 probability that two local unitary corrections are needed. Moreover, the two single qubit measurements are in different basis, observing that the physical implementation of (matter) qubit measurement is usually only native in one basis, we may consider that an additional single-qubit unitary is needed to transform measurement basis. Therefore, to generate $N$-qubit GHZ state in quantum networks based on gate teleportation: $(3N-2)$ qubits are needed, $(2N-2)$ of which are dedicated to $(N-1)$ Bell pairs; $(2N-2)$ single-qubit measurements are needed; $(2N-2)$ local CNOTs are needed; and on average $(N-1)$ (or $(2N-2)$) single-qubit gates are needed. The estimation is summarized in Table. 2.
VI.2 GHZ merging
Merging of two GHZ states can be achieved with local CNOT and measurement feedforward, as shown explicitly in the following where we consider the merging of two GHZ states with $N$ and $M$ qubits, respectively:
$|\mathrm{GHZ}_{N}\rangle|\mathrm{GHZ}_{M}\rangle\propto(|\underbrace{0\dots 0}_{N\times~{}0}\rangle+|\underbrace{1\dots 1}_{N\times~{}1}\rangle)(|\underbrace{0\dots 0}_{M\times~{}0}\rangle+|\underbrace{1\dots 1}_{M\times~{}1})\rangle$ (112)
${\mathrm{CNOT}_{N,N+1}}$ $(|\underbrace{0\dots 0}_{(N-1)\times~{}0}00\underbrace{0\dots 0}_{(M-1)\times~{}0}\rangle+|\underbrace{0\dots 0}_{(N-1)\times~{}0}01\underbrace{1\dots 1}_{(M-1)\times~{}1}\rangle+|\underbrace{1\dots 1}_{(N-1)\times~{}1}11\underbrace{0\dots 0}_{(M-1)\times~{}0}\rangle+|\underbrace{1\dots 1}_{(N-1)\times~{}1}10\underbrace{1\dots 1}_{(M-1)\times~{}1}\rangle).$
Then it is obvious that if we measure the target qubit of CNOT (with index $N+1$ when all qubits are indexed from 1 through $N+M$) in computational basis, when the measurement outcome is 0, the post-measurement state is exactly a GHZ state with $N+M-1$ qubits, and when the outcome is 1, the post-measurement state can be transformed into a GHZ state with $N+M-1$ qubits by applying $\min[N,M-1]$ local X gates to flip the $|0\rangle$ and $|1\rangle$. An example circuit of GHZ merging is visualized in Fig. 10(b).
We can also estimate the resources needed for this approaches. Different from using gate teleportation, merging does not consume all qubits in generated Bell pairs. We need to perform $(N-1)$ mergings, and for each merging there is 1/2 probability that one local unitary correction is needed, and 1/2 probability that no local unitary correction is needed. Therefore, to generate $N$-qubit GHZ state in quantum networks based on merging: $(2N-2)$ qubits are needed, all of which are dedicated to $(N-1)$ Bell pairs; $(N-1)$ single-qubit measurements are needed; $(N-1)$ local CNOTs are needed; and on average $(N-1)/2$ single-qubit gates are needed. The estimation is summarized in Table. 2.
However, consider the separation between communication qubits and sensor qubits, we need to first generate the GHZ states across communication qubits from merging and then perform local SWAP gates to swap the GHZ state from communication qubits to the sensor qubits. In this case, we need $N$ sensor qubits, and $N$ local SWAP gate between communication and sensor qubits (which can be decomposed into 3 CNOTs).
Table 2: Resource estimation for $N$-qubit GHZ state assembly from $(N-1)$ Bell pairs by CNOT teleportation and GHZ merging.
qubits [^1] Including $(2N-2)$ qubits of the Bell pairs. 1-qubit measurements 2-qubit gates 1-qubit gates [^2] Not including measurement basis transformation.
CNOT teleportation $23N-$ 2 $22N-$ 2 $22N-$ 2333All are CNOTs. $N-1$
[…]
VII Network simulation details
In this section we provide details of how we simulate the initial probe state preparation for DQS using SeQUeNCe.
VII.1 Entanglement distribution
The preparation of multipartite entangled probe state starts with the distribution of bipartite entangled states between pairs of network nodes. These bipartite entangle states will be assembled into multipartite entangled states using local operations and classical communication, such as gate teleportation and GHZ state merging (graph state fusion).
The distribution of bipartite entanglement involves remote entanglement generation between nearest network nodes that are directly connected by a physical implementation of quantum channels (e.g., optical fibers or free-space optics), entanglement swapping which extends the generated entangled states to more distant node pairs, and potentially entanglement purification which is expected to improve the quality (fidelity) of distributed entangled pairs.
For entanglement generation, in SeQUeNCe we implement a new abstract model of single-heralded entanglement generation protocol based on meet-in-the-middle photonic Bell state measurement. The underlying processes assumed for a single attempt of entanglement generation includes:
  1. 1.
    Local memory-photon entanglement on a network node is generated, and the photon is transmitted to a middle interference center via a lossy optical fiber to perform heralded measurement (Bell state measurement/BSM).
  2. 2.
    Two photons being transmitted from both nodes are expected to simultaneously arrive at the middle node at certain time determined by fiber lengths.
  3. 3.
    However, due to the existence of losses in optical fibers, it is possible that one or both photons are not successfully transmitted. We assume that the heralded measurement can only be successful if both photons arrive.
  4. 4.
    Moreover, we assume that the underlying implementation of heralded measurement consists of linear optics only, which fundamentally upper bounds the success probability of BSM to 1/2 [130].
According to the above description, the overall success probability of an entanglement generation attempt is $p_{t,l}p_{t,r}p_{m}$ where $p_{t,l(r)}$ is the probability for left (right) photon to arrive at the middle station and $p_{m}$ is the success probability of heralded measurement when both photons arrive, which in simulation is a positive, tunable parameter below 1/2. Additionally, the successfully generated entangled state, is assumed to be a Bell diagonal state (BDS).
Entanglement swapping involves three nodes, one middle node which will perform projective BSM, and two end nodes which will receive the BSM results from the middle node and perform local corrections accordingly. We consider the standard quantum circuit of BSM as shown in Fig. 11(a), where measurements are both in computational basis. We also consider that different physical implementation of entanglement swapping could lead to different success probability, and allow the success probability of each entanglement swapping to be tunable. Bipartite entanglement purification involves two nodes which share multiple entangled state. In SeQUeNCe, the default protocol is the BBPSSW/DEJMPS protocol based on bilocal CNOT, i.e. both nodes will perform CNOT on two qubits from two entangled states they hold, and perform single qubit computational basis measurement on one qubit per node, whose circuit is shown in Fig. 11(b). The measurement outcomes from both nodes are classically communicated to each other to determine if the purification attempt is successful.
Figure 11: Quantum circuits for (a) standard Bell state measurement (BSM), and (b) entanglement purification protocol based on bilocal CNOT, specifically the BBPSSW/DEJMPS protocol.
To increase the practicality in our simulation, we consider imperfections in single-qubit measurements and two-qubit gates (especially CNOT). However, we do not consider errors in single-qubit gates, because experimentally multi-qubit gates are much more noisy than single-qubit ones, with infidelity about one order of magnitude higher. As we assume entangled states are in BDS form, we perform offline analytical derivation of imperfect entanglement swapping results in Sec. VII.5.1, and imperfect entanglement purification results in Sec. VII.5.2. These allow us to avoid explicit tracking of density matrix under quantum operations during simulation of entanglement distribution. We assume a homogeneous model of gate and measurement imperfections on each node: On one specific node, the error model of multi-qubit gate is unchanged when applying to different qubits, and so is the error model of single-qubit measurement.
We note that when multiple entangled states (quantum memories) are available, the space of strategies for executing entanglement protocols (e.g., generation, swapping and purification) is large. Currently in SeQUeNCe, the strategy is determined by the priorities of each protocol in the stack. For concreteness, our implementation gives the highest priority to entanglement purification (i.e., purification will be performed ASAP), whenever multiple entangled states are available on a single link (between two network nodes). Entanglement swapping has the second highest priority, and it will be performed immediately when there is one entangled state established between a center node and each of its left and right neighbors (not necessarily nearest neighbors). We also emphasize that when taking into account quantum memory decoherence, and the non-universality of entanglement purification [131], the optimization of quantum protocol strategies/policies [132] will be very hard due to the large policy space, so this is beyond the scope of this work where simulation is for the demonstration of principle for DQS probe state preparation.
VII.2 Time-dependent quantum memory decoherence
Quantum memories inevitably undergo decoherence throughout the time after they are initialized. In quantum networks where communication takes longer time than local quantum information processing due to long distance between network nodes, the effect of memory idling decoherence can be even more significant. Therefore, besides imperfections in quantum operations such as gate and measurement, in this work we also implement time-dependent memory decoherence which is analytically modeled by continuous time Pauli channels [133] which is naturally compatible with the BDS assumption, to reflect the noise effect of storing entangled states in quantum memories before final usage. Note that besides idling decoherence, quantum states, especially entangled states distributed over the network, are changed only when quantum protocols are operated, such as entanglement swapping and entanglement purification. Therefore, memory idling decoherence effect only needs to be added to the quantum state prior to the quantum protocols, according to the duration of idling which can be calculated as the difference between the current time and the last time point when the quantum state is updated. In this way, the quantum memory decoherence over time smoothly fits in the discrete event simulation framework.
Due to the decoherence over time, one cannot use a certain quantum memory forever, and resetting is needed after a certain amount of time, which is the cutoff time for the storage of quantum states. In our simulation, we set the cutoff time to be proportional to quantum memory coherence time, while the coefficient is a tunable parameter. Specifically, for entangled states involving more than one quantum memories, the time for all involved memories to be reset is determined by the smallest reset time among the memories.
We also have additional comment on entanglement generation. Note that the quantum memories which store the successfully generated entangled states are initialized before the nodes receive the heralding signals. To account for the idling errors of quantum memories before the arrival of measurement results, we assume that the successfully generated entangled state starts from a certain BDS, specified by the entanglement generation protocol (parameters including the initial fidelity and relative strength of three Pauli error components). Then the entangled state will decohere under independent quantum memory decoherence channels which are dependent on idling time, and the decoherence time for each memory is determined by the time when the memory is initialized and the time when the heralding signal is received.
VII.3 Probe state preparation as network application
Our objective probe state is a global GHZ state of all sensors distributed in different spatial locations. The generation of this GHZ state requires entangling operation which is realized by distributed Bell pairs by the quantum network. Conceptually, we may view the process of generating global GHZ state as an application of the quantum network, which consumes distributed bipartite entangled states from the service of quantum network. More specifically, the application can be divided into two parts:
  1. 1.
    The involved $N$ sensor nodes in the network collectively establish this application, and decide one out of the $N$ nodes as the center node which corresponds to the center qubit of standard GHZ state preparation circuit. Then the rest $(N-1)$ nodes request bipartite entanglement, i.e. Bell pair, with the center node. This is standard bipartite entanglement distribution in quantum repeater network.
  2. 2.
    When bipartite entanglement has been established between the sensor nodes, it can be used to perform CNOT teleportation, or GHZ merging.
At the end of network simulation, the assembly process of all the distributed Bell pairs into a GHZ state is implemented with the help of functions from the open-source package QuTiP [134]. Specifically, we provide two implementations of the process, one using GHZ merging and the other using CNOT teleportation, where we explicitly simulate noisy CNOT gates and noisy single-qubit measurement, together with the classical feedforward correction based on single-qubit measurement outcome(s). Specifically, in this work we implement the assemble-in-the-end protocol, where bipartite entanglement between sensor nodes are requested within a fixed period of time, before they are assembled into a multipartite state. We note that the optimization of the assembly process, e.g. optimal time of assembly in analogy to optimal time of purification [133], can be interesting future work. It is then possible that more than one entangled link between two sensor nodes exist at the end the stage, and in such cases we perform a final round of purification before the assembly, so that between each pair of nodes there is at most one entangled link. The final purification is implemented in a fidelity-aware manner (assuming the capability of estimating entangled state fidelities based on information of system hardware and timing), due to the consideration of the non-universality of entanglement purification [131]: We always attempt to purify the two lowest-fidelity entanglement pairs in the entangled state ensemble, and we repeat this process until there are at most one entangled pair left. After this process, we utilize QuTiP functions to simulate the assembly of bipartite entangled states into a GHZ state.
We re-emphasize that the processes during bipartite entanglement distribution can be easily tracked with SeQUeNCe features, including noisy entanglement generation, swapping and purification, and time dynamics for noisy Bell state in BDS form under quantum memory idling channel. Moreover, the network simulation with SeQUeNCe is parallelizable [135], so in principle we can simulate bipartite entanglement distribution at large scale, which is physically allowed because in such scenarios entangled states are independent bipartite states. However, when we start to create multipartite entanglement, the necessity of storing larger quantum states and even quantum dynamics simulation will eventually limit the problem scale, especially for DQS where multipartite entangled resource state is desired.
VII.4 Network simulation parameters
Here we enumerator the key tunable system parameters of network simulation in this work.
VII.4.1 Protocol configuration
For the probe state generation protocol, there is mainly one parameter, the cutoff time $T_{\mathrm{pc}}$: The GHZ state generation protocol consists of generation of individual bipartite entanglement, and the final assembly of the GHZ state. All pairs of sensor nodes will attempt Bell state generation for $T_{\mathrm{pc}}$, before the final assembly.
VII.4.2 Network configuration
  1. 1.
    Network topology $\mathcal{G}$. It is a graph that determines how the nodes, including both sensor nodes themselves and potential intermediate repeater nodes in the network core, are directly connected with physical channels, e.g. optical fibers.
  2. 2.
    Channel length $L$. It is the length of a specific optical fiber, which will determine the loss of photons.
  3. 3.
    Classical communication time $T_{\mathrm{cc}}$. We consider that classical communication may require some higher level communication protocols, thus the time is not necessarily equal to length divided by the speed of light.
  4. 4.
    Number of quantum memories per node $M$. This represents how many quantum memories are available for bipartite entanglement generation attempts.
VII.4.3 Hardware configuration
  1. 1.
    Quantum memory coherence time $\tau_{m}$. It determines the time scale of how quantum states stored in quantum memories degrade over idling.
  2. 2.
    Quantum memory error pattern. It determines the ratio between three Pauli errors in the continuous time idling channel model.
  3. 3.
    Quantum memory frequency $f_{m}$. It determines how frequently the entanglement generation can be attempted.
  4. 4.
    Quantum memory efficiency $\eta_{m}$. It determines the probability of successfully establishing memory-photon entanglement in each entanglement generation attempt.
  5. 5.
    Memory cutoff ratio $r_{m}$. After a certain quantum memory has been initialized for $r_{m}\tau_{m}$ time, it will be reset.
  6. 6.
    Raw Bell state $\rho_{0}$. It is a BDS that is the initial state for entanglement generation, before any quantum memory decoherence occurs.
  7. 7.
    Two-qubit gate fidelity $\eta_{\mathrm{g}}$ and single-qubit measurement fidelity $\eta_{\mathrm{m}}$. They determine the performance of quantum operations in the quantum network, including entanglement swapping, entanglement purification, and GHZ merging or gate teleportation.
VII.5 Analytical modeling
Here we provide the analytical formulae which facilitate our simulation. They could also be of independent interest to other analytical studies. In the following, for both entanglement swapping and purification, we include imperfection in gates (we assume that single-qubit gate imperfections is negligible w.r.t. multi-qubit, especially two-qubit in our setup, gates and thus do not account them in the following) and measurements [71] involved in the teleportation protocol: $\tilde{U}_{ij}\rho\tilde{U}^{\dagger}_{ij}=pU_{ij}\rho U^{\dagger}_{ij}+\frac{1-p}{4}I_{ij}\otimes\mathrm{tr}_{ij}\rho$ and $\tilde{P}_{i=0,1}=\eta|i\rangle\langle i|+(1-\eta)|1-i\rangle\langle 1-i|$, where $\mathrm{tr}_{ij}(\cdot)$ represents partial tracing over qubits $i,j$, $U_{ij}$ is an ideal two-qubit unitary, and $\tilde{U}_{ij}$ is an imperfect implementation of $U_{ij}$, which has $p$ probability of perfect implementation and $(1-p)$ probability of resulting in depolarizing error. $\tilde{P}_{i=0,1}$ is the POVM corresponding to imperfect implementation of single-qubit projective measurement $P_{i}=|i\rangle\langle i|$, which has $\eta$ probability of giving a correct measurement outcome. We note that some of the results derived below have been used under specific settings in [136] written by part of the research team of this work, and the special case of identical measurement and gate error rates for two parties, i.e. $\eta_{1}=\eta_{2}=\eta$ and $p_{1}=p_{2}=p$, can be found in [137] without derivation. In the future we may consider other operation error models, which could make analytical derivation of results more complicated and thus explicit simulation of quantum operations [138] during network simulation might then be necessary.
VII.5.1 Entanglement swapping
Consider two Bell diagonal states, one between node A and node B and the other between node B and node C, where node B is the middle swapping station:
$\rho_{\mathrm{A,B_{1}}}(\vec{\lambda})\otimes\rho_{\mathrm{B_{2},C}}(\vec{\lambda^{\prime}})=\left(\lambda_{1}\Phi^{+}+\lambda_{2}\Phi^{-}+\lambda_{3}\Psi^{+}+\lambda_{4}\Psi^{-}\right)_{\mathrm{A,B_{1}}}\otimes\left(\lambda_{1}^{\prime}\Phi^{+}+\lambda_{2}^{\prime}\Phi^{-}+\lambda_{3}^{\prime}\Psi^{+}+\lambda_{4}^{\prime}\Psi^{-}\right)_{\mathrm{B_{2},C}},$ (113)
where $|\Phi^{\pm}\rangle=(|00\rangle\pm|11\rangle)/\sqrt{2}$, $|\Psi^{\pm}\rangle=(|01\rangle\pm|10\rangle)/\sqrt{2}$, and $\Phi^{\pm},\Psi^{\pm}$ are the projectors onto the corresponding pure Bell states. Entanglement swapping requires performing BSM on the two qubits held by node B, i.e. $\mathrm{B_{1},B_{2}}$. Conditioned on measurement outcome, single-qubit operation is further performed on qubit A (or C) to ensure that the resulting state is in a specific form of Bell state, and here we focus on $\Phi^{+}$ without loss of generality. Specifically, if the outcomes (in the order of $\mathrm{B_{1},B_{2}}$) are 00 do nothing; if the outcomes are 01, apply $X$ gate; if the outcomes are 10, apply $Z$ gate; if the outcomes are 11, apply $Y$ gate.
Through straightforward evaluation of CNOT on $\mathrm{B_{1},B_{2}}$ + Hadamard on $\mathrm{B_{1}}$ for tensor products of two pure Bell states, and the four imperfect POVM’s corresponding to four possible measurement outcomes, we can obtain that the final output state is:
$\rho_{A,C}(\vec{\lambda},\vec{\lambda^{\prime}})=$ $p[\eta_{1}\eta_{2}C_{I}+(1-\eta_{1})\eta_{2}C_{X}+\eta_{1}(1-\eta_{2})C_{Z}+(1-\eta_{1})(1-\eta_{2})C_{Y}]\Phi^{+}_{\mathrm{A,C}}$
$+p[(1-\eta_{1})\eta_{2}C_{I}+\eta_{1}\eta_{2}C_{X}+(1-\eta_{1})(1-\eta_{2})C_{Z}+\eta_{1}(1-\eta_{2})C_{Y}]\Phi^{-}_{\mathrm{A,C}}$
$+p[\eta_{1}(1-\eta_{2})C_{I}+(1-\eta_{1})(1-\eta_{2})C_{X}+\eta_{1}\eta_{2}C_{Z}+(1-\eta_{1})\eta_{2}C_{Y}]\Psi^{+}_{\mathrm{A,C}}$
$+p[(1-\eta_{1})(1-\eta_{2})C_{I}+\eta_{1}(1-\eta_{2})C_{X}+(1-\eta_{1})\eta_{2}C_{Z}+\eta_{1}\eta_{2}C_{Y}]\Psi^{-}_{\mathrm{A,C}}$
$+\frac{1-p}{4}(\Phi^{+}_{\mathrm{A,C}}+\Phi^{-}_{\mathrm{A,C}}+\Psi^{+}_{\mathrm{A,C}}+\Psi^{-}_{\mathrm{A,C}}),$ (114)
where we have defined $C_{I}=\lambda_{1}\lambda_{1}^{\prime}+\lambda_{2}\lambda_{2}^{\prime}+\lambda_{3}\lambda_{3}^{\prime}+\lambda_{4}\lambda_{4}^{\prime},\ C_{X}=\lambda_{1}\lambda_{2}^{\prime}+\lambda_{2}\lambda_{1}^{\prime}+\lambda_{3}\lambda_{4}^{\prime}+\lambda_{4}\lambda_{3}^{\prime},\ C_{Y}=\lambda_{1}\lambda_{4}^{\prime}+\lambda_{4}\lambda_{1}^{\prime}+\lambda_{2}\lambda_{3}^{\prime}+\lambda_{3}\lambda_{2}^{\prime},\ C_{Z}=\lambda_{1}\lambda_{3}^{\prime}+\lambda_{3}\lambda_{1}^{\prime}+\lambda_{2}\lambda_{4}^{\prime}+\lambda_{4}\lambda_{2}^{\prime}$, and we assume CNOT error probability $(1-p)$ and single-qubit measurement error probabilities $(1-\eta_{1})$ and $(1-\eta_{2})$.
VII.5.2 Entanglement purification
Under these error models, consider that CNOT gates on both sides have different error probabilities $p_{A}$ and $p_{B}$, also measurements have different error probabilities $\eta_{A}$ and $\eta_{B}$. Then for two Bell diagonal states as input to the DEJMPS purification protocol, the (un-normalized) output state conditioned on success is:
$\tilde{\rho}_{A_{1},B_{1}}(\vec{\lambda},\vec{\lambda^{\prime}})=$ $p_{A}p_{B}\left[(1-\eta_{A}-\eta_{B}+2\eta_{A}\eta_{B})(\lambda_{1}\lambda_{1}^{\prime}+\lambda_{2}\lambda_{2}^{\prime})+(\eta_{A}+\eta_{B}-2\eta_{A}\eta_{B})(\lambda_{1}\lambda_{3}^{\prime}+\lambda_{2}\lambda_{4}^{\prime})\right]\Phi^{+}_{\mathrm{A_{1},B_{1}}}$
$+p_{A}p_{B}\left[(1-\eta_{A}-\eta_{B}+2\eta_{A}\eta_{B})(\lambda_{1}\lambda_{2}^{\prime}+\lambda_{2}\lambda_{1}^{\prime})+(\eta_{A}+\eta_{B}-2\eta_{A}\eta_{B})(\lambda_{1}\lambda_{4}^{\prime}+\lambda_{2}\lambda_{3}^{\prime})\right]\Phi^{-}_{\mathrm{A_{1},B_{1}}}$
$+p_{A}p_{B}\left[(1-\eta_{A}-\eta_{B}+2\eta_{A}\eta_{B})(\lambda_{3}\lambda_{3}^{\prime}+\lambda_{4}\lambda_{4}^{\prime})+(\eta_{A}+\eta_{B}-2\eta_{A}\eta_{B})(\lambda_{3}\lambda_{1}^{\prime}+\lambda_{4}\lambda_{2}^{\prime})\right]\Psi^{+}_{\mathrm{A_{1},B_{1}}}$
$+p_{A}p_{B}\left[(1-\eta_{A}-\eta_{B}+2\eta_{A}\eta_{B})(\lambda_{3}\lambda_{4}^{\prime}+\lambda_{4}\lambda_{3}^{\prime})+(\eta_{A}+\eta_{B}-2\eta_{A}\eta_{B})(\lambda_{3}\lambda_{2}^{\prime}+\lambda_{4}\lambda_{1}^{\prime})\right]\Psi^{-}_{\mathrm{A_{1},B_{1}}}$
$+\frac{1-p_{A}p_{B}}{8}(\Phi^{+}_{\mathrm{A_{1},B_{1}}}+\Phi^{-}_{\mathrm{A_{1},B_{1}}}+\Psi^{+}_{\mathrm{A_{1},B_{1}}}+\Psi^{-}_{\mathrm{A_{1},B_{1}}}),$ (115)
where we use $A_{1},B_{1}$ to denote the two qubits of kept entangled pair and $A_{2},B_{2}$ have been measured (traced out), while the BDS density matrix elements with prime correspond to the measured entangled pair. Then the success probability is just the trace of the above un-normalized state
$p_{s}=$ $p_{A}p_{B}[\eta_{A}\eta_{B}+(1-\eta_{A})(1-\eta_{B})](\lambda_{1}\lambda_{1}^{\prime}+\lambda_{2}\lambda_{2}^{\prime}+\lambda_{1}\lambda_{2}^{\prime}+\lambda_{2}\lambda_{1}^{\prime}+\lambda_{3}\lambda_{3}^{\prime}+\lambda_{4}\lambda_{4}^{\prime}+\lambda_{3}\lambda_{4}^{\prime}+\lambda_{4}\lambda_{3}^{\prime})$
$+p_{A}p_{B}[\eta_{A}(1-\eta_{B})+(1-\eta_{A})\eta_{B}](\lambda_{1}\lambda_{3}^{\prime}+\lambda_{2}\lambda_{4}^{\prime}+\lambda_{1}\lambda_{4}^{\prime}+\lambda_{2}\lambda_{3}^{\prime}+\lambda_{3}\lambda_{1}^{\prime}+\lambda_{4}\lambda_{2}^{\prime}+\lambda_{3}\lambda_{2}^{\prime}+\lambda_{4}\lambda_{1}^{\prime})$
$+\frac{1-p_{A}p_{B}}{2},$ (116)
from which we can explicitly obtain normalized output state conditioned upon success as $\rho_{A_{1},B_{1}}(\vec{\lambda},\vec{\lambda^{\prime}})=\tilde{\rho}_{A_{1},B_{1}}(\vec{\lambda},\vec{\lambda^{\prime}})/p_{s}$. Specifically, the first diagonal element in Bell basis corresponding to $\Phi^{+}_{\mathrm{A_{1},B_{1}}}$ is the fidelity
$F_{s}=\frac{p_{A}p_{B}\left[\frac{\eta_{A}\eta_{B}+(1-\eta_{A})(1-\eta_{B})}{2}(\lambda_{1}\lambda_{1}^{\prime}+\lambda_{2}\lambda_{2}^{\prime})+\frac{\eta_{A}(1-\eta_{B})+(1-\eta_{A})\eta_{B}}{2}(\lambda_{1}\lambda_{3}^{\prime}+\lambda_{2}\lambda_{4}^{\prime})\right]+\frac{1-p_{A}p_{B}}{16}}{p_{s}/2}.$ (117)
Additionally, we can prove that no matter what input (Bell diagonal) states are, no matter what CNOT infidelities and measurement infidelities are, the probability of getting measurement outcome indicating success is always not lower than 1/2:
Proposition VII.1.
Every trial of CNOT based recurrence entanglement protocol which has Bell diagonal states as input, using imperfect CNOT and single qubit measurement whose error models are described above, will get measurement outcome indicating success with probability not lower than 1/2.
Proof.
We use Equation 116 as the starting point. Notice that $\lambda_{1}+\lambda_{2}+\lambda_{3}+\lambda_{4}=\lambda_{1}^{\prime}+\lambda_{2}^{\prime}+\lambda_{3}^{\prime}+\lambda_{4}^{\prime}=1$ according to normalization and $21\geq\lambda_{1},\lambda_{1}^{\prime}\geq 1/$ 2 to ensure that the BDS’s are entangled. After some reorganization we have
$p_{s}=$ $\frac{1}{2}+p_{A}p_{B}[\eta_{A}(1-\eta_{B})+(1-\eta_{A})\eta_{B}]$
$+p_{A}p_{B}[ab+(1-a)(1-b)][\eta_{A}\eta_{B}+(1-\eta_{A})(1-\eta_{B})-\eta_{A}(1-\eta_{B})-(1-\eta_{A})\eta_{B}]-\frac{1}{2}p_{A}p_{B},$ (118)
where we have defined $a\coloneqq\lambda_{1}+\lambda_{2}$, $b\coloneqq\lambda_{1}^{\prime}+\lambda_{2}^{\prime}$, and thus naturally $21\geq a,b\geq 1/$ 2. Then we have
$p_{s}\geq$ $\frac{1}{2}+p_{A}p_{B}\left([\eta_{A}(1-\eta_{B})+(1-\eta_{A})\eta_{B}]+\frac{1}{2}[\eta_{A}\eta_{B}+(1-\eta_{A})(1-\eta_{B})-\eta_{A}(1-\eta_{B})-(1-\eta_{A})\eta_{B}]-\frac{1}{2}\right)$
$=$ $\frac{1}{2}+p_{A}p_{B}\left(\frac{1}{2}[\eta_{A}\eta_{B}+(1-\eta_{A})(1-\eta_{B})+\eta_{A}(1-\eta_{B})+(1-\eta_{A})\eta_{B}]-\frac{1}{2}\right)$
$=$ $\frac{1}{2}+p_{A}p_{B}\left(\frac{1}{2}-\frac{1}{2}\right)=\frac{1}{2},$ (119)
where for the first inequality we have used the fact that $ab+(1-a)(1-b)\geq 1/2$ for $21\geq a,b\geq 1/$ 2. ∎